A Complete Bibliography of Publications in the ACM Symposia on Theory of Computing

(STOC) for 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.09

Title word cross-reference

#CSP [DR10].

1 [Bel12]. 2 [CST13, GKL12]. 2^{log^{1−ϵ} n} [KPV12]. 3 [ASSS12]. 4 [GKL12]. = [JJUW10]. Δ [ASS13]. ϵ [SA12]. F_{p}^{n} [HL11]. h [DHiK11]. K [IP11, ASSS12, BCC^{+}10, COP12, LS13, Mol12, MS11, Tho13]. k^{2} [BDF^{+}11]. L_{1} [LS11, SW11]. l_{1}^{n} [Kar11]. l_{p} [Kar11]. n/ log(n) [VV11]. n_{2+o(1)} [Ull13]. O(m) [OSV12]. o(n) [CS13a]. o(n^{1/2}) [iKK11]. O(n^{1/4}) [BCC^{+}10]. o(n log n) [GKK10]. O(nm) [Orl13]. p < 2 [Kar11]. R^{3} [HJN12]. s [AKS12]. t [AKS12].

- [AKS12]. -approximation [iKK11, SA12]. -certificates [Bel12]. -colourings [Mol12]. -median [LS13, IP11]. -minor-free [DHiK11].

-NAESAT [COP12]. -norm [SW11]. -sample [VV11]. -SAT [MS11]. -subgraph [BCC^{+}10]. -systems [CST13]. -time [OSV12].

’13 [ACM13]. 17th [BC10].

2006 [ACM06].

abstract [Ajt12, BNT13, Bel12, Ber13, CGW13, HKL^{+}11, Kay12, KS13, KNP11, LW11, LPS13, NSV11]. accepting [Hua13]. ACM [ACM06, ACM10, ACM11, ACM12, ACM13]. adaptive [HW13]. advantage [Amb13]. adversaries [VV12]. adversary [Ajt11]. advice [AD10]. AdWords [GML12]. Affine [Kay12, BFH^{+}13, HL11, ZBS11]. affine-invariant [BFH^{+}13]. after [COP13]. against [BGJK12, DV10, VV12]. agreement [KS13, MH13]. algebraic [BSGK^{+}13, GX13]. algebraic-geometric [GX13]. algebraization [LN10]. algorithm [AGMS11, AKS12, BCY11, CFJ^{+}10, Chu11, FHZ11, GMSV12, HPR13, iKW10, iKW11, KOSZ13, MV10, MS11, NSV11, OSV12, SA12, Vég12]. algorithmic [ALSV13, CDW12, DHiK11, HL10]. algorithms [Amb13, CMSS13, CCEL12, CKL12, EK13, ESY12, FGR^{+}13, GTW13, HKL^{+}11, IKL^{+}11, INSWN11, iKK11, KLL^{+}13, Mad10, MMV12, Mon10, NS13, Orl10, RW13, Yos11]. allocating [KO11]. allows [DvM10]. Almost [ACER11, CHSS11, CLP10, FFGL13]. alternating [JNS13]. Alto [ACM13]. amplification [BHP10, CKOR10, Li12]. amplifying [HN12]. analysis [BIKK12, BR12, CP12, HR13, KC10, KLL^{+}13]. analyst [HRU13]. analysts [MO11]. Analyzing [Hae11]. ancestry [FK10]. Annual [ACM06, ACM13]. Answering [Ull13]. any [Kar11, LP11]. applications [AMS12, ALSV13, BDYW11, BBH^{+}12, BP13, BH13b, CC13, CKL12, CPS13, DHiK11, GGSW13, KM13, MM13, SW11]. approach [BGT10, BM13, LRS13, MY11]. Approximate [AdFM11, GLPS10, ACG12, ALSV13, BP13, KMP12, MH13, NTZ13, ZBS11]. approximately [KM11]. Approximating [LS13, OSV12, FL11]. Approximation [BHM10, CCEL12, Cha13, Hua13, MMV12, BGK12, BCC^{+}10, BGRS10, CKM^{+}11, CW13, DDFS12, iKK11, LS13, LY13, Mad10, RW13, SHK^{+}11, SA12, Woo11, Yos11]. Approximations [RST10, BH13a]. arbitrary [FG10, GHK^{+}12]. arguments [GW11]. arithmetic [Ajt12, Raz10]. arrivals [MY11]. assumptions [Ajt10, ABW10, GW11, Pas11]. asymptotically [CKOR10]. asynchronous [MH13]. Attribute [GVW13]. Attribute-based [GVW13]. auctions [BGGM10, Dob11, DFK11, DRY11, FFGL13, GML12, LR12, PP11]. Augmenting [Vég10]. automata [BGJ13]. Average [KR13b, DHK^{+}10b]. Average-case [KR13b]. averages [SS13].

balanced [OSV12]. balancing [GW10, LW11]. balloons [CCVV12]. barrier [BDF^{+}11, GHRU11]. base [DPT10]. based [BSGK^{+}13, BGRS10, DK11, GVW13, IP11, MY11, MV10]. bases [CKN13]. basing [AGGM06, AGGM10]. Bayesian [BCGL12, HL10]. be [RK11]. Beating [HR12]. better [Orl13]. Beyond [HR13, CCD13]. bidders [DFK11, LR12]. Bilipschitz [LM10a]. bimatrix [AGMS11]. bin [ACKS13]. binomial [DDS12]. bipartite [GKK10, KMT11, MY11, SA12]. Black [SV11, CIL12, CPS13, Goy13, Woo11, AFW12]. Black-box [SV11, CIL12]. Blackbox [SS11]. blanket [DLP11]. block [ACER12]. Boolean [CS13a]. bootstrapping [BCCT13]. Bottom [Tho13]. Bottom- [Tho13]. bound [CR11, FGR^{+}13, GW12, GX13, VZ10]. bounded [BHM10, GHK^{+}12, GTW13, Hir10, KRR13, KMSV10, SS11, Yos11]. bounded-degree [Yos11]. bounded-depth [GHK^{+}12]. Bounding [DHK^{+}10b]. bounds [ACER11, ASSS12, Ajt13, ACKS13, BDYW11, BBI12, BDL13, BKS12, BI13b, CS13b, CP12, CLP10, FMP^{+}12, FHZ11, GHK^{+}12, KR13b, LS11, LW11, NN13, Pat10, Raz10, She10, She13, Wil10, WZ12]. box [CIL12, CPS13, Goy13, SV11, Woo11]. BQP [Aar10]. branching [DMPY12, ESY12]. Breaking [BDF^{+}11, iKK11]. Budget [BCGL12, BGGM10]. buffer [ACER11, ACER12]. buffering [VZ10]. build [AN12]. Byzantine [KS13, MH13].

CA [ACM11]. calculus [BNT13]. California [ACM13]. Cambridge [ACM10]. can [RK11]. canonical [BW13]. capacitated [LMM13]. capacity [KR13a]. capture [CGW13]. carrying [BCCT13]. case [HR13, KR13b]. cases [NTZ13]. Catching [COP12]. cell [Lar12, MV10]. Certifiable [VV12]. certificates [Bel12]. Changing [DPT10]. channel [KR13a]. characterization [AD10, CDW12, CFW12]. characterized [BFH^{+}13]. Characterizing [VZ12]. checkable [BSCGT13]. checking [CKN13]. Cheeger [KLL^{+}13, LGT12]. chow [DDFS12]. Christofides [AKS12]. cipher [HKT11]. circuit [PP10]. circuits [ASSS12, GHK^{+}12, GKP^{+}13, GVW13, GKL12, HWY10, KMSV10, MV13, SV11, SS11]. class [Vég12]. Classical [BLP^{+}13, RK11]. Claus [Sve11]. clearing [Orl10]. clinching [GML12]. cliques [FGR^{+}13]. clocks [BNS13]. closest [KPV12]. cloud [LATV12]. CLTs [VV11]. clustering [FL11, IP11]. codes [BDYW11, BSGK^{+}13, Efr12, GHK^{+}12, GHK10, GX12, KL12, KS10, KSY11, Woo13]. coding [BR11, Hae11]. Coevolutionary [BGM13]. collapses [DvM10]. colourings [Mol12]. Combinatorial [BGT10, FGL13, BDYW11, Dob11, DRY11, GMRZ11, KOSZ13, KLP12, LY13]. commitments [LP11]. Communication [She13, BBCR10, BR11, BGPW13, CR11, HN12, KLPP11, RK11, She11, She12b]. commutative [HWY10]. Competitive [GK12, GGK13]. complementary [GMSV12]. complete [AMS12, BGJ13, CGW13, CFW12]. completion [JNS13]. complex [CC12]. Complexity [CC12, KC10, AA11, BHP10, BCZ13, Bra12, BM13, BI11, BI13b, CR11, CPY13, DV12, DR10, GJ10, HN12, Kop11, Lar12, NSV11, PP10, She11, She12b, SS13, TZ13]. Composable [ST13]. composition [BCCT13]. compress [BBCR10]. compressed [FS12]. compressive [IP11]. computation [Ajt11, Ajt12, BGJK12, CHHKM12, GHW11, GJ10, HR13, HRU13, KLO10, LATV12, RVW13]. computational [AA11, DV12]. computations [MV10]. Computing [ACM06, ACM10, ACM11, ACM12, ACM13, AGKM12, GHK^{+}12, LRS13, Orl10, SS13]. concave [DJ12, GMSV12]. concrete [BSCGT13]. concurrent [CCEL12, Goy13]. condensers [Li12]. condition [CFW12]. Conditional [Sve10]. conditioned [TS13]. conductance [CHHKM12, CLP10]. congestion [Chu12b, iKK11]. conjecture [HJN12, RS10]. conjunctions [GHRU11]. conjunctive [Mar10]. connected [CHHKM12]. Connectivity [DP10, Vég10]. Constant [Goy11, LP11, Bel12, Chu12b, GGK13, Yos11]. constant-competitive [GGK13]. Constant-round [LP11]. constant-sized [Bel12]. constant-time [Yos11]. constrained [BGGM10, Sve10]. Constraint [HS13, BK12, Mar10]. constructing [HRV10]. construction [Kar11, LPS13]. constructions [VZ12]. contagion [GK12]. contention [VCZ11]. context [ESY12]. context-free [ESY12]. contingency [KRSU10]. continual [BGJK12, DNPR10]. continuous [CKM^{+}13]. Contraction [DHiK11]. contracts [CCVV12]. convergence [Smi11]. convex [DRY11, Vég12]. coordination [CCG^{+}11]. Coppersmith [Wil12]. correctable [BDYW11, BSGK^{+}13]. correcting [GHK^{+}12]. correlated [DFK11, KRSU10]. Correlation [HL11]. cost [Vég12]. counter [BGJ13]. counting [CC12, Lar12, MN12, Ull13]. coupled [KMP12]. Cover [DLP11, OR10, Var10]. covert [GJ10]. credit [KO11]. crossing [Chu11]. cryptographic [Ajt10]. cryptography [ABW10, BP13]. CSP [CC12, Yos11]. CSPs [TZ13]. cubic [HS10]. cut [BNS13, CFW12, GTW13, INSWN11, KM13]. cuts [LRTV12]. cutset [MR11]. cycle [iKR10]. cycles [DHK10a].

D [ASSS12]. dark [ACG13]. data [BCCT13, FL11, KNPW11]. datasets [BO10a]. decidable [GGRÀ10]. decodability [GHK10, Woo13]. decodable [Efr12]. decoding [GX12, GX13, KS10, KSY11]. decomposable [AMS12]. decomposition [DHiK11, iKW11]. decompositions [AN12, CC13, KMS13]. deferral [GGK13]. degree [ASSS12, BSGK^{+}13, Yos11]. degree- [ASSS12]. degree-lifted [BSGK^{+}13]. delegating [RVW13]. Delegation [KRR13]. deletions [Ber13]. dense [KKM13]. densest [BCC^{+}10]. densities [BCC^{+}10]. dependence [CHHKM12]. depth [ASSS12, ASS13, GHK^{+}12, GKL12, KMSV10, SV11, SS11]. depth- [ASSS12, GKL12]. depth-3 [SS11]. depth-4 [KMSV10, SV11]. depth-D [ASSS12]. derandomization [MS11, Wil13]. derivatives [She13]. descent [CCD13]. Design [Li12, BDYW11, BCGL12, CHMS10, CIL12, CST13, FT11, HL10]. designs [BW13]. Detecting [BCC^{+}10, FGR^{+}13]. determinant [AS10, CHSS11, HT12]. Determinism [Ajt12]. Deterministic [KMSV10, Kar11, BGJ13, MV10, NS13]. devices [ACER12]. diameter [RW13]. dice [VV12]. dichotomy [CGW13]. different [ABW10]. Differential [DNPR10, HRU13, HT10, NTZ13, Ull13]. differentially [BDKT12]. dimension [GS11]. dimensional [BIKK12, CDW12, Kar11]. dimensionality [BGK12, NN13]. dimensions [COW13]. direct [Kla10, She11]. Directed [DK11, Ber13]. directional [She13]. discrepancy [MN12]. discrete [BL13, DMN13]. disjoint [iKK11]. disjointness [Kla10, She12b]. distance [ACG12, CR11, HPR13, IP11]. distances [AGMP13]. distortion [FHS11, LS11, MM13]. Distributed [KLO10, SHK^{+}11, GHW11, WZ12]. distributions [BL13, DDS12, KMT11]. divine [FT11]. do [GHW11]. doesn’t [SS11]. Don’t [PT11a]. doubling [LS11]. duality [ZBS11]. Dueling [IKL^{+}11]. dynamic [ACG12, KLO10, Lar12, Mad10, NS13, Pat10, VZ10].

earth [IP11]. easy [DFK11, Hae11]. Edge [KL12, iKK11]. edge-disjoint [iKK11]. edit [AGMP13]. Efficiency [HRV10, BSCGT13]. Efficient [NRV13, FHS11, FFGL13, ST13]. Efficiently [KMV10]. eigenvalues [LRTV12]. Eighth [ACM06]. Electrical [CKM^{+}11, LRS13]. elimination [Ajt13]. ellipsoid [CFJ^{+}10]. embedding [LS11]. embeddings [FHS11, MM13, SW11]. encryption [GGSW13, GKP^{+}13, GVW13, LATV12]. enough [CFW12]. entangled [KV11]. entropy [CKOR10, VV11]. equations [BC10, KM11]. equilibrium [HRU13, FGL13]. Equivalence [BGJ13, HKT11]. err [FT11]. Erratum [AGGM10]. error [BCZ13, GHK^{+}12, HL11, KS10, Woo13]. error-correcting [GHK^{+}12]. errors [BLP^{+}13, BR11]. Estimating [VV11]. estimation [KNPW11, Smi11]. estimator [VV11]. ETH [BI13a]. Euclidean [ES13, HPR13]. evasive [DL12]. Every [BFH^{+}13, NS11, Yos11]. Exact [HKL^{+}11, Amb13, BGPW13]. excluded [GM12]. exclusion [BBJK12, GW12]. exhaustive [Wil10]. existence [KLP12]. expansion [BY12, RS10]. expected [KS13]. Explicit [BI13b, BDF^{+}11, FHS11]. explicitly [Tal10]. exponential [BNS13, FMP^{+}12, Li13, MV10, OSV12]. exponentially [RK11]. expressions [MMN10]. extended [Ajt12, BNT13, Bel12, Ber13, BM13, CGW13, FMP^{+}12, HKL^{+}11, Kay12, KS13, KNP11, LW11, LPS13, NSV11]. Extending [CKM^{+}13]. Extensions [LM10b]. external [VZ10]. extractors [DV10, Li12, Li13, ZBS11].

factor [MY11]. factor-revealing [MY11]. factorization [AGKM12]. failure [DP10]. falsifiable [GW11]. families [BDL13]. family [BSGK^{+}13]. fan [GKL12, KMSV10]. fan-in [GKL12, KMSV10]. fanin [SS11]. Fast [CKL12, CKN13, KNPW11, LPS13, RW13, CHHKM12]. Faster [KMP12, Mad10, CKM^{+}11, Wil12]. feasible [BCGL12]. few [Hua13]. Fibonacci [BLT12]. field [GX12, SS11]. fields [Kop11]. fifth [ACM13]. find [PT11a]. Finding [CCVV12, GiKMW11]. Finetti [BH13b]. fingerprints [AGMP13]. finite [Kop11, TZ13]. finite-valued [TZ13]. Fisher [Orl10]. Fixed [MR11, GiKMW11]. Fixed-parameter [MR11, GiKMW11]. flow [CCEL12, CKM^{+}11, DK11, EK13, INSWN11, KMP12, Mad10, Vég12]. flow-based [DK11]. flows [CKM^{+}11, KMP12, LRS13, Orl13]. fly [LATV12]. Folded [GX12]. forbidden [ACG12]. forbidden-set [ACG12]. forest [BHM10]. forgive [FT11]. form [BW13]. formation [BGM13]. formula [KR13b]. formulas [ASSS12, ASS13, DMPY12, Raz10]. formulations [BM13, FMP^{+}12]. Forty [ACM13]. Forty-fifth [ACM13]. Fourier [CP12, HIKP12]. FPT [BDT11]. fractional [Mad10]. fractionality [Hir10]. framework [FL11, FHHP11]. free [BCGL12, DHiK11, ESY12, LR12]. freeness [iKY13]. freezing [Mol12]. frequency [BO10b]. full [AD10, MS11]. Fully [ACG12, Goy13, LATV12, NS13]. function [GX12, LP11, VCZ11]. functional [GKP^{+}13, WZ12]. functions [AGGM06, AGGM10, App12, BH11, Bel12, CS13a, CPS13, DHK^{+}10b, Goy11, HRV10, KM13, MZ10].

Gabidulin [GX13]. games [AGMS11, BGM13, HKL^{+}11, KV11, OW12, RS10]. gap [CR11, KLL^{+}13]. gap-Hamming-distance [CR11]. gaps [GORZ12]. garbled [GKP^{+}13]. gates [GHK^{+}12]. Gaussians [KMV10]. general [FHHP11]. generation [VV12]. generator [VZ12]. generators [App12, GMRZ11, HRV10, KNP11, MZ10]. Geometric [BI11, BI13b, GX13, SA12, Var10]. geometry [BDYW11, BSGK^{+}13, HT10, NTZ13]. Gibbard [MR12]. Global [CHHKM12]. Going [COP13]. good [KL12]. gossip [Hae11]. gradient [CCD13]. grammars [ESY12]. Graph [RS10, CC13, Chu11, FHHP11, iKW10, iKW11, iKY13, Mad10, Mil13, Mol12]. graphs [ACG12, AMS12, BHM10, BGT10, Ber13, BP11, CKM^{+}11, Chu12b, DHiK11, DP10, EK13, GKK10, GM12, GTW13, INSWN11, KL12, KMS13, NS11, RST10, RW13]. gross [CCD13]. Grothendieck [NRV13]. ground [BH13a]. group [KNP11]. groups [MV13].

halfspace [MN12]. halfspaces [DDFS12, She10]. Hamiltonicity [CKN13]. Hamming [CR11]. hard [Ull13]. Hardness [BHP10, AGGM06, AS10, BLP^{+}13, CCEL12, CHSS11, GTW13, HN12, KM11, KPV12, OW12, SHK^{+}11, Sve10, AGGM10]. hashing [PT11b]. heaps [BLT12]. heterogeneous [BGGM10]. hidden [AC12]. hierarchy [Aar10, DvM10]. High [KSY11, BCC^{+}10, HL11, Kar11, KS10]. High-rate [KSY11]. higher [COW13, KLL^{+}13, LGT12, LRTV12]. higher-order [LGT12]. hits [ASSS12]. hitting [ASSS12, ASS13]. hitting-set [ASS13]. hitting-sets [ASSS12]. holds [BI13a]. HOM [GGRÀ10]. homeomorphism [AGMS11]. homologous [DHK10a]. Homomorphic [AGMP13, LATV12]. human [FT11]. Hypercontractivity [BBH^{+}12]. hypercube [CS13a]. hypercubes [CS13b]. hyperfinite [NS11]. hypergraph [Mar10]. hypergraphic [GORZ12]. hypergraphs [GW10, KKM13]. hypergrids [CS13b]. ideal [HKT11]. identical [Sve10]. identities [HT12]. identity [FS12, KMSV10, SV11, SS11]. implementations [GHW11]. implies [BGK12, Wil10]. impossibility [BP13, Dob11]. Improved [BR12, INSWN11, KLL^{+}13, Orl10, BGRS10, GX12]. improvement [Li13]. improvements [HRV10]. Improving [AKS12, Wil10]. inapproximability [Yos11]. incoherent [HR12]. independence [BO10a, Tho13]. independent [Cha13, CHMS13, Li13]. individual [CCVV12]. induced [AMS12]. inequalities [KW12, LGT12]. inequality [KLL^{+}13, NRV13]. information [Ajt11, Bra12, BM13, BGPW13, CFJ^{+}10, FHS11]. Inner [CCG^{+}11]. input [CW13, MM13]. input-sparsity [MM13]. inputs [FT11, HW13, Hua13]. instances [Hua13]. integral [CCEL12]. integrality [GORZ12]. Interactive [Bra12, KR13a, RR10, RVW13, BBCR10, BR11, GW11]. International [ACM10, ACM11, ACM12]. interpolation [BGT10]. intersection [She10]. invariance [HKM10]. invariant [BFH^{+}13, HL11]. Inverting [TS13]. irreducible [Efr12]. isomorphism [CST13, GM12]. isoperimetric [RST10]. item [PP11]. items [BGGM10].

Jacobian [ASSS12]. jobs [Sve11]. Johnson [DKS10]. Jose [ACM11]. June [ACM10, ACM11, ACM13].

k-SAT [COP13]. key [ABW10, LLW11].

labeling [BKS12]. labels [ACG12]. Lanczos [OSV12]. lanky [ES13]. Laplacian [CKM^{+}11]. Large [CC13, AMS12, Mil13, OR10, Woo13]. Large-treewidth [CC13]. lattice [CMSS13, MV10]. laws [BO10b]. LDPC [KL12]. leak [LLW11]. leakage [BGJK12]. leaking [Ajt11]. Learning [BH11, DDS12, BLP^{+}13, KMV10]. Lee [SS13]. lemma [AKS10, HS13]. lifted [BSGK^{+}13]. limit [AFW12]. Limits [Pas11, BGT10, CIL12, LM10b, VZ10]. Lindenstrauss [DKS10]. Linear [EK13, FMP^{+}12, AA11, BDKT12, DHK10a, DK11, GMSV12, GHK10, HPR13, HW13, KOSZ13, KM11, KMS13, KS10, MM13, NSV11, SA12, Woo13]. Linear-time [EK13]. Linearizable [GHW11]. linkage [iKW10]. Lipschitz [CS13b, KM13]. List [GX13, GHK10, GX12, KS10, Woo13]. list-decodability [GHK10]. list-decoding [KS10]. LLL [NSV11]. LLL-reduction [NSV11]. Load [GW10, LW11]. Local [KS10, AKS10, App12, BH13b, HS13, LSV10]. locality [App12]. locally [BDYW11, BSGK^{+}13, BFH^{+}13, Efr12]. locking [FHS11]. log [BCC^{+}10]. log-densities [BCC^{+}10]. logspace [TS13]. long [App12]. looking [Mon10]. losing [DPT10]. loss [ACG13, CKOR10]. Lovasz [HS13, AKS10, KS11]. Low [CW13, JNS13, MM13, AN12, App12, BGK12, DDFS12, FHS11, FS12]. low-dimensionality [BGK12]. Low-distortion [MM13, FHS11]. Low-rank [JNS13, FS12]. low-weight [DDFS12]. Lower [Ajt13, ASSS12, BBI12, BKS12, BI13b, CR11, FGR^{+}13, FMP^{+}12, FHZ11, GW12, KR13b, NN13, Pat10, Raz10, She13, VZ10, Wil10]. LP [BGRS10].

LP-based [BGRS10]. LPs [MY11].

MA [ACM10]. machines [Sve10, Sve11]. made [Hae11]. Maintaining [Ber13, OR10, GGK13]. Majority [DMN13]. majorizing [DLP11]. Making [She12a]. malleable [Goy11, Li12, LP11]. management [ACER11]. Many [LRTV12, Tal10]. maps [CKM^{+}13, NN13]. market [Orl10]. markets [CPY13, GMSV12]. matching [BDL13, DJ12, KMT11, LSV10, MY11, NS13, OR10, SA12]. matchings [AMS12, CKN13, GKK10, KKM13]. matrices [BDYW11, BDF^{+}11, HR12, Kan10, KRSU10, TS13, Wil12]. matrix [ALSV13, AGKM12, CKL12, JNS13]. Matroid [KW12, LSV10]. Matroids [GORZ12]. matter [SS11]. matters [GS11]. max [EK13, INSWN11, Orl13]. maximal [NS13]. maximization [VCZ11]. maximum [BR11, CKM^{+}11, Hir10, LRS13]. May [ACM06, ACM12]. measurements [BH13b, GLPS10]. measures [DLP11]. Measuring [BO10a]. Mechanism [FT11, BCGL12, CHMS10, CIL12, HL10, RR10]. Mechanisms [KO11, BDKT12, CDW12, CHMS13, CCG^{+}11, DRY11, ST13]. median [IP11, LS13, RR10]. meet [KS11]. meets [AFW12, iKY13]. membership [AdFM11, VZ10]. memory [BGJK12, VZ10]. Message [Mon10]. messages [LPS13]. method [BGT10, OSV12]. methods [Kan10]. metrics [LM10a]. min [INSWN11]. minimal [Tho13]. Minimax [AFW12]. minimization [JNS13]. minimum [Vég12]. minimum-cost [Vég12]. minor [DHiK11, iKW10, iKW11]. MinSum [CCG^{+}11]. mis [KO11]. misalignments [AGMP13]. mixtures [KMV10]. model [HKT11, IP11, MMN10, VZ10]. model-based [IP11]. moment [KNPW11]. money [AC12]. monitoring [WZ12]. Monotone [BY12, CP12, CPY13]. monotonicity [CS13a, CS13b]. Moser [KS11]. most [MV10]. mover [IP11]. Multi [CHMS10, CST13, LGT12, CDW12, ESY12]. multi-dimensional [CDW12]. Multi-parameter [CHMS10]. Multi-stage [CST13]. multi-type [ESY12]. Multi-way [LGT12]. multicommodity [KMP12, Mad10]. Multicut [BDT11, MR11]. Multidimensional [MH13]. multiflow [CFW12, Hir10]. multikey [LATV12]. multilinear [DMPY12, GKL12, KMSV10, SV11, VCZ11]. multiobjective [BR12]. Multiparty [BGJK12, LATV12, She12b]. multiple [EK13]. multiple-source [EK13]. Multiplying [Wil12]. multiway [BNS13]. mutual [BBJK12, GW12].

NAESAT [COP12]. Natural [Wil13]. Near [DV10, LS11, Woo11, BC10, SA12]. near-linear [SA12]. Near-optimal [DV10, LS11, Woo11]. Nearly [AMS12, DDFS12, HIKP12, KOSZ13, KMST10]. nearly-linear [KOSZ13]. negative [LM10a]. Net [HPR13]. network [Hae11]. networks [CFW12, CP12, DFF11, FG10, GK12, KLO10]. neutrality [MR12]. NL [BGJ13]. NL-complete [BGJ13]. no [CHHKM12, DvM10]. node [LMM13, Vég10]. node-capacitated [LMM13]. node-connectivity [Vég10]. nodes [Chu12a]. noise [DHK^{+}10b, She12a]. Non [CPS13, Goy13, HWY10, CPY13, GW11, Goy11, Li12, LP11]. Non-black-box [CPS13, Goy13]. Non-commutative [HWY10]. non-interactive [GW11]. non-malleable [Goy11, Li12, LP11]. non-monotone [CPY13]. noncommutative [AS10, CHSS11, NRV13]. nondeterminism [Ajt12]. nonnegative [AGKM12]. nontrivial [DvM10]. norm [FHS11, SW11]. NP [AGGM06, KM11, OW12, AGGM10]. NP-hardness [AGGM06, KM11, OW12, AGGM10]. number [Chu11, VV12]. NY [ACM12].

obfuscation [BP13]. objectives [Vég12]. Oblivious [Ajt10]. observation [DNPR10]. occur [ASSS12]. occur- [ASSS12]. Odd [iKR10]. off [BNT13]. offs [HN12]. Okamura [LMM13]. On-the-fly [LATV12]. one [AGGM06, AGGM10, App12, BGJ13, BIKK12, BO10b, CPS13, Goy11, HRV10, LP11, RK11, Vég10]. one-counter [BGJ13]. one-dimensional [BIKK12]. one-way [AGGM06, AGGM10, App12, CPS13, HRV10, LP11, RK11]. Online [DJ12, KMT11, MY11, ACER12, ACKS13, BKS12, GGK13]. operators [KC10]. opinion [BGM13]. optics [AA11]. Optimal [ACER12, CS13b, DHK10a, DFK11, ES13, GS11, MN12, She10, Yos11, CR11, CKOR10, DV10, DDFS12, DRY11, FK10, GX12, HIKP12, KNPW11, KMST10, LS11, MO11, PP11, Smi11, VV11, Woo11]. optimization [BR12, DRY11, LY13, Mil13]. optimizing [GLPS10]. option [AFW12]. oracle [HKT11]. oracles [ACG12, DP10]. orbit [COW13]. order [KLL^{+}13, LGT12]. ordered [LR12]. orientability [GW10].

packet [HS13]. packing [ACKS13, iKR10]. pairwise [Cha13]. Palo [ACM13]. Parallel [KV11, CFW12, LW11]. parameter [CHMS10, GiKMW11, MR11]. parameterized [MR11]. parameters [DDFS12, RST10]. parenthesized [MMN10]. Pareto [MO11]. partial [CFJ^{+}10]. partitioning [BNS13, KLL^{+}13, LGT12, MMV12]. passing [Mon10]. path [AKS12, GS11]. paths [Ber13, EK13, iKK11]. Perfect [GKK10, CKN13, KKM13]. petal [AN12]. petal-decompositions [AN12]. piecewise [GMSV12]. piecewise-linear [GMSV12]. pivot [GMSV12]. pivoting [FHZ11]. planar [ACG12, BHM10, EK13, INSWN11, KMS13]. planted [FGR^{+}13]. point [OW12]. Poisson [DDS12, LY13]. Polyhedral [GML12]. Polynomial [ESY12, KKM13, Aar10, AGMS11, ASS13, BGK12, BNT13, BC10, DvM10, DHK^{+}10b, KS13, MZ10, Pat10, Vég12]. Polynomial-time [KKM13, DvM10]. polynomiality [CKM^{+}13]. polynomials [HS10, KM13, Kay12, She10, She12a]. polytope [AdFM11, GML12]. polytopes [HKM10]. poorly [CHHKM12]. posets [FK10]. posted [CHMS10]. power [GGK13, LSV10, PT11b]. powering [Kop11]. precedence [Sve10]. predicates [Hua13]. preprocessing [KPV12]. preserving [Smi11]. PRG [KM13]. price [KRSU10]. prices [Orl10]. pricing [AFW12, CHMS10]. principle [HKM10]. Prior [CHMS13, LR12, BCGL12]. Prior-free [LR12, BCGL12]. Prior-independent [CHMS13]. priority [Tho13]. Privacy [CKOR10, Smi11, DNPR10, HT10, HRU13, Li12, NTZ13, RR10, Ull13]. Privacy-preserving [Smi11]. private [BDKT12, HR13, HRU13, MN12, Woo11]. Privately [GHRU11, KRSU10]. Probabilistic [KLP12]. probabilistically [BSCGT13]. probabilistically-checkable [BSCGT13]. probe [Lar12]. problem [BGK12, BCY11, BNS13, BKS12, BC10, COW13, Chu11, DDFS12, GGRÀ10, HWY10, iKK11, KPV12]. problems [BK12, CFW12, HPR13, Hir10, Mad10, MMV12, MV10, Mil13, Pat10, Vég12]. Proceedings [ACM06, ACM10, ACM11, ACM12, ACM13]. processes [ESY12]. Product [BH13a, CCG^{+}11, Kla10, She11]. Product-state [BH13a]. products [KNP11]. profile [RST10]. programming [DHK10a]. programs [Bel12, DK11, DMPY12]. projections [Kay12]. prone [DP10]. proof [BHP10, BCCT13, HN12, iKW10, iKW11]. proof-carrying [BCCT13]. proofs [AM12, BBH^{+}12, BSCGT13, HT12, HN12, RVW13, Wil13]. propeller [HJN12]. properties [HL11, Mar10]. property [BFH^{+}13, iKY13, NS11]. prophet [KW12]. protocols [Goy11, Woo11]. provable [Pas11]. provably [AGKM12]. proximity [RVW13]. prune [HPR13]. pseudo [LS13]. pseudo-approximation [LS13]. pseudoentropy [VZ12]. Pseudorandom [App12, GMRZ11, KNP11, MZ10, HRV10, VZ12]. PSPACE [JJUW10]. Public [ABW10]. Public-key [ABW10].

QIP [JJUW10]. quadratically [KMP12]. quantifier [Ajt13]. quantitative [MR12]. Quantum [AC12, BH13b, RK11, AD10, AKS10, Amb13, BCY11, BH13a, DV10, She11, TS13, VV12]. quartic [HS10]. Quasi [ASS13, NSV11, Var10]. quasi-linear [NSV11]. Quasi-polynomial [ASS13]. quasi-uniform [Var10]. Quasipolynomial [BW13, BCY11, CST13]. Quasipolynomial-time [BW13, BCY11, CST13]. queries [AdFM11, BDKT12, Mar10, Ull13]. query [DV12, GHRU11, She11].

radius [RW13]. Ramanujan [KL12]. RAMs [Ajt13, Ajt10]. Random [CMSS13, App12, BGT10, GW10, GHK10, HKT11, KRSU10, KS10, MY11, MMV12, Mol12, VV12, Woo13]. randomized [DRY11, FHZ11, GW12, GHW11, HR12, LW11]. range [Lar12]. Rank [AGMS11, BDYW11, ALSV13, BI11, CKL12, CW13, FS12, JNS13, Raz10]. Rank-1 [AGMS11]. rate [GX12, KSY11]. rates [Smi11, Woo13]. Rational [AM12]. really [ES13]. reals [KM11]. Recognizing [MMN10]. Reconstruction [GKL12]. recovery [FS12, GLPS10, IP11]. Recursive [BCCT13, KMS13]. red [CCVV12]. reducing [NN13]. reduction [NSV11]. reductions [CIL12]. Reed [GX13]. regime [HL11]. regression [CW13, MM13]. regular [BI13a, GKK10]. related [RST10]. relations [FHS11]. relaxation [VCZ11]. relaxations [GORZ12]. releasing [GHRU11, KRSU10]. reordering [ACER11]. repetition [KV11]. representations [Efr12]. representing [She10]. rescue [CCD13]. resettable [BP13, CPS13]. resistance [Cha13, Hua13]. resolution [BBI12, BI13a, VCZ11]. response [HR12]. result [Dob11]. results [BNT13, GTW13, Yos11]. returns [DJ12]. Reusable [GKP^{+}13]. revealing [MY11]. revisited [HKT11]. rigid [KLP12]. RIP [BDF^{+}11]. rises [CGW13]. RMR [GW12]. Robust [BK12, HW13, MM13, She12a]. robustness [CCVV12]. roots [PT11a]. round [GJ10, Goy11, LP11]. rounding [NRV13]. Routing [Chu12b, HS13, LPS13]. rows [KRSU10]. rules [FHZ11]. rumor [CHHKM12]. rumors [DFF11]. rumour [CLP10]. rush [PT11a].

salesman [BGK12]. sample [VV11]. sampling [BL13, Tho13, Var10]. San [ACM11]. Santa [Sve11]. SAT [COP13, MS11]. satisfaction [BK12, HS13, Mar10]. Satisfiability [DvM10, BK12, PP10]. satisfiable [Hua13]. Satterthwaite [MR12]. Saving [LN10]. scaling [BGT10]. Schaefer [BP11]. schedules [Sve11]. scheduling [ACER12, CHMS13, Sve10]. Schelling [BIKK12]. scheme [BGK12, FK10]. schemes [BHM10, Mad10, VCZ11]. Scholes [AFW12]. Schöning [MS11]. scientific [KO11]. SDD [KOSZ13]. search [GS11, LSV10, Wil10]. searchability [FG10]. Seattle [ACM06]. section [Kar11]. Secure [Ajt11, BGJK12, VV12]. security [CPS13, Pas11]. segregation [BIKK12]. selfishness [CCVV12]. semi [MMV12]. semi-random [MMV12]. semidefinite [FMP^{+}12]. sensing [FS12, IP11]. sensitivity [DHK^{+}10b]. separability [BCY11]. separable [GMSV12, Vég12]. Separating [DMPY12, GW11]. separation [FMP^{+}12]. separator [KMS13, OSV12]. sequential [CHMS10]. series [CFW12]. series-parallel [CFW12]. serving [ACG13]. set [ACG12, ASS13, She12b, Tho13, Var10]. set-depth- [ASS13]. sets [ASSS12, DL12, Tal10]. setting [Goy13]. settling [CHSS11]. Seymour [LMM13]. shapes [GMRZ11]. Shielding [MV13]. shift [AGMP13]. Short [HT12, ES13]. shorter [iKW10, iKW11]. shortest [Ber13, EK13]. shown [VV11]. sign [She10]. sign-representing [She10]. signatures [CGW13]. similarity [Tho13]. Simple [NS13, KOSZ13, PT11b]. simpler [iKW11]. Simplex [BNS13, FHZ11]. simplifying [VZ12]. simulation [CPS13, Goy13]. Simultaneous [FFGL13]. single [MV10, PP11]. single-item [PP11]. singleton [GX13]. singular [HR13]. size [KR13b, MR11, VV11]. sized [Bel12]. sketches [HW13]. sketching [AGMP13]. Smale [BC10]. small [FK10, FG10, GS11, LPS13, OR10, Tal10]. small-world [FG10]. smoothed [BR12, BC10, MO11]. SNARKS [BCCT13]. snowflakes [LM10a]. Social [DFF11]. Solomon [GX13]. Solution [HJN12, BC10]. solutions [DDFS12, MO11]. Solving [BC10, Mil13, HKL^{+}11, KOSZ13, KM11]. Some [BNT13]. Sorting [CFJ^{+}10]. SoS [DMN13]. source [EK13, Li13, ZBS11]. space [BBJK12, BBI12, DPT10, HN12, KRR13, KNPW11, LN10]. spaces [CCG^{+}11, LS11]. Span [Bel12]. spanners [DK11, ES13]. spanning [AN12]. sparse [BGT10, DKS10, GLPS10, HIKP12, IP11, LRTV12, NTZ13, RW13]. Sparsest [GTW13, KM13]. sparsification [DvM10, FHHP11, KMST10, LM10b]. sparsifiers [Chu12a]. Sparsity [NN13, CW13, MM13]. spectra [KRSU10]. Spectral [Kan10, KLL^{+}13, LGT12, Mil13, OSV12, RST10]. split [CCVV12]. spread [DFF11]. spreading [CHHKM12, CLP10]. squares [BBH^{+}12, HWY10]. stablest [DMN13]. stage [CST13]. standard [Pas11]. state [BH13a]. states [BH13a]. Statistical [FGR^{+}13, GHRU11, Smi11]. Steiner [BW13, BHM10, BGRS10, CST13, Chu12a, GORZ12, GGK13]. STOC [ACM13]. STOC’06 [ACM06]. STOC’10 [ACM10]. STOC’11 [ACM11]. STOC’12 [ACM12]. Stochastic [LY13, ESY12, HKL^{+}11]. storage [DV10]. streaming [MMN10]. streams [KNPW11]. stretch [AN12, App12]. Strict [BLT12]. Strong [BI13a, She11, FMP^{+}12, Kla10]. stronger [RK11]. Strongly [Vég12, MY11]. structural [iKY13]. Structure [GM12, CMSS13, FG10, HS10]. Structured [KMS13]. structures [KLP12]. subcodes [GX13]. subdivision [iKY13]. subdivision-freeness [iKY13]. Subexponential [FHZ11]. Subgraph [KMST10, BCC^{+}10]. subgraphs [GiKMW11, GM12]. subgroups [Cha13]. sublinear [KSY11, RVW13]. sublinear-time [KSY11]. sublogarithmic [DFF11]. Submodular [VCZ11, BH11, Dob11]. subset [Tho13]. Subspace [DL12, SW11, MM13]. subspaces [AC12]. substitutes [CCD13]. success [Mon10]. Succinct [BL13, GW11, GKP^{+}13, HN12]. suffice [GHW11]. sum [BBH^{+}12, HWY10]. sum-of-squares [BBH^{+}12, HWY10]. sums [Tho13]. Superlinear [Amb13, BBI12]. superpolynomial [BBI12, Wil10]. support [VV11]. switching [CP12]. symmetric [KL12]. Symposium [ACM06, ACM10, ACM11, ACM12, ACM13]. systems [CST13, CKM^{+}11, KOSZ13, MH13].

table [LPS13]. tables [KRSU10]. tabulation [PT11b]. take [PT11a]. Tardos [KS11]. Tatonnement [CCD13]. Tensor [Raz10, BI11]. Tensor-rank [Raz10]. tensors [FS12, Kan10]. test [GM12]. testable [BFH^{+}13, NS11]. tester [CS13a]. Testing [iKY13, CS13b, CST13, FS12, HL11, KMSV10, KS10, SV11, SS11]. tests [Ajt12]. their [AMS12, BBH^{+}12]. theorem [BP11, GM12, iKW10, Kla10, LMM13, MR12]. theorems [BH13b, She11, SS13]. theoreticians [Mon10]. Theory [ACM06, ACM10, ACM11, ACM12, ACM13, BI11, BI13b, KC10, iKY13, Mil13]. thin [ES13]. Thirty [ACM06]. Thirty-Eighth [ACM06]. threshold [COP13, COP12, DHK^{+}10b, MZ10, Mol12]. thresholds [GW10]. Tight [ACKS13, BBJK12, BKS12, CP12, GHK^{+}12, LW11, WZ12, ACER11, CLP10, GW12, VZ10]. Time [BBI12, AGMS11, BW13, BBJK12, BGK12, BCY11, BC10, CST13, CW13, DvM10, DFF11, EK13, ESY12, GLPS10, GKK10, HPR13, HN12, KKM13, KOSZ13, KS13, KMS13, KSY11, MM13, MV10, NSV11, OSV12, Orl13, PT11a, RVW13, SA12, Yos11]. Time-space [BBI12, BBJK12, HN12]. times [DLP11]. top [GKL12, KMSV10, SS11]. topological [GiKMW11, GM12]. topology [KLPP11]. total [DHK10a]. towers [GX12]. tractability [MR11]. Tractable [Mar10, GiKMW11]. trade [BNT13, HN12]. trade-off [BNT13]. trade-offs [HN12]. tradeoff [BBJK12]. tradeoffs [BBI12]. transcendence [ASSS12]. transform [DKS10, HIKP12]. transformation [Woo11]. transitive [KL12]. traveling [BGK12]. tree [AN12, BGRS10, GORZ12, GGK13]. treewidth [BHM10, CC13, GTW13]. trial [BCZ13]. triangulations [CMSS13]. true [VV12]. truthful [Dob11]. TSP [AKS12]. two [KMV10, iKK11, She10, ZBS11]. two-source [ZBS11]. type [ESY12, LM10a].

ultrasparsifiers [KMST10]. uncertain [FT11]. uncertainty [FHS11]. Unconditional [BDKT12, Yos11]. undecidability [CKM^{+}13]. underlying [FG10]. undirected [CKM^{+}11, Chu12b, INSWN11, Vég10]. unified [FL11]. uniform [Var10]. unimodularity [DHK10a]. union [PT11a]. unique [iKW10, OW12, RS10]. unit [EK13]. unit-weight [EK13]. universal [FK10]. unknown [KMT11]. unless [DvM10]. unrelated [Sve11]. unseen [VV11]. updates [LLW11]. USA [ACM06, ACM10, ACM11, ACM12, ACM13]. Using [AN12, Goy11, JNS13, KMP12, LRS13, LPS13, Mil13, She13]. utilities [GMSV12].

valuations [Dob11]. valued [TZ13]. vanishing [CGW13]. vector [ACKS13, BDL13, HR13, KPV12]. verification [SHK^{+}11]. versus [Ajt12, Wil13]. vertex [Chu12a, LM10b, OR10]. via [ACG12, BNS13, BI13b, CP12, CKN13, DK11, HRU13, LS13, LY13, LATV12, LRTV12, Mad10, MN12, RR10, VV11, Var10, VCZ11, Woo11, ZBS11]. virtue [HN12]. Voronoi [MV10]. vs [FMP^{+}12].

WA [ACM06]. Walrasian [FGL13]. way [AGGM06, AGGM10, App12, CPS13, Goy11, HRV10, LGT12, LP11, RK11]. weight [DDFS12, EK13]. Weighted [Var10, Ber13]. weights [CC12]. well [MMN10, TS13]. well-parenthesized [MMN10]. Winograd [Wil12]. wireless [KLPP11]. without [Ajt10, CFJ^{+}10, DPT10, MR12]. Witness [GGSW13]. world [CHHKM12, FG10]. worlds [GS11]. worst [HR13]. worst-case [HR13].

Yang [SS13]. York [ACM12].

Zero [BO10b]. Zero-one [BO10b].

References

Aaronson:2011:CCL

[AA11] Scott Aaronson and Alex Arkhipov. The computational complexity of linear optics. In ACM [ACM11], pages 333–342. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Aaronson:2010:BPH

[Aar10] Scott Aaronson. BQP and the polynomial hierarchy. In ACM [ACM10], pages 141–150. ISBN 1-60558-817-2. LCCN QA 76.6

.A152 2010. URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Applebaum:2010:PKC

[ABW10] Benny Applebaum, Boaz Barak, and Avi Wigderson. Publickey cryptography from different assumptions. In ACM [ACM10], pages 171–180. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010.

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Aaronson:2012:QMH

[AC12] Scott Aaronson and Paul Christiano. Quantum money from hidden subspaces. In ACM [ACM12], pages 41–60. ISBN 1-4503-

1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Adamaszek:2011:ATB

[ACER11] Anna Adamaszek, Artur Czumaj, Matthias Englert, and Harald Räcke. Almost tight bounds for reordering buffer management. In ACM [ACM11], pages 607–616. ISBN 1-4503-0691-8. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Adamaszek:2012:OOB

[ACER12] Anna Adamaszek, Artur Czumaj, Matthias Englert, and Harald Räcke. Optimal online buffer scheduling for block devices. In ACM [ACM12], pages 589–598. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Abraham:2012:FDA

[ACG12] Ittai Abraham, Shiri Chechik, and Cyril Gavoille. Fully dynamic approximate distance oracles for planar graphs via forbidden-set distance labels. In ACM [ACM12], pages 1199–1218. ISBN 1-

4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Azar:2013:LSD

[ACG13] Yossi Azar, Ilan Reuven Cohen, and Iftah Gamzu. The loss of serving in the dark. In ACM [ACM13], pages 951–960. ISBN 1-4503-2029-5.

Azar:2013:TBO

[ACKS13] Yossi Azar, Ilan Reuven Cohen, Seny Kamara, and Bruce Shepherd. Tight bounds for online vector bin packing. In ACM [ACM13], pages 961–970. ISBN 1-4503-2029-5.

ACM:2006:SPT

[ACM06] ACM, editor. STOC’06: Proceedings of the Thirty-Eighth Annual ACM Symposium on Theory of Computing 2006, Seattle, WA, USA, May 21–23, 2006. ACM Press, New York, NY, USA, 2006. ISBN 1-59593-134-1. LCCN QA75.5 .A22 2006. URL http://

portal.acm.org/citation.cfm?id=1132516. ACM order num-

ber 508060.

ACM:2010:SPA

[ACM10] ACM, editor. STOC’10: Proceedings of the 2010 ACM International Symposium on Theory of Computing: June 5–8, 2010, Cambridge, MA, USA. ACM Press, New York, NY, USA, 2010. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

ACM:2011:SPA

[ACM11] ACM, editor. STOC’11: Proceedings of the 2011 ACM International Symposium on Theory of Computing: June 6–8, 2011, San Jose, CA, USA. ACM Press, New York, NY, USA, 2011. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

ACM:2012:SPA

[ACM12] ACM, editor. STOC’12: Proceedings of the 2012 ACM International Symposium on Theory of Computing: May 19–22, 2012, New York, NY, USA. ACM Press, New York, NY, USA, 2012. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

ACM:2013:SPF

[ACM13] ACM, editor. STOC ’13: Proceedings of the Forty-fifth Annual ACM Symposium on Theory of Computing: June 1–4, 2013, Palo Alto, California, USA. ACM Press, New York, NY, USA, 2013. ISBN 1-4503-2029-5.

Aaronson:2010:FCQ

[AD10] Scott Aaronson and Andrew Drucker. A full characterization of quantum advice. In ACM [ACM10], pages 131–140. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Arya:2011:APM

[AdFM11] Sunil Arya, Guilherme D. da Fonseca, and David M. Mount. Approximate polytope membership queries. In ACM [ACM11], pages 579–586. ISBN 1-4503-0691-8. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Abernethy:2012:MOP

[AFW12] Jacob Abernethy, Rafael M. Frongillo, and Andre Wibisono. Minimax option pricing meets Black–Scholes in the limit. In ACM [ACM12], pages 1029–1040. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Akavia:2006:BOW

[AGGM06] Adi Akavia, Oded Goldreich, Shafi Goldwasser, and Dana Moshkovitz. On basing one-way functions on NP-hardness. In ACM [ACM06], pages 701–710. ISBN 1-59593-134-1. LCCN

QA75.5 .A22 2006. URL http://portal.acm.org/citation.

cfm?id=1132516. See erratum [AGGM10].

Akavia:2010:EBO

[AGGM10] Adi Akavia, Oded Goldreich, Shafi Goldwasser, and Dana Moshkovitz. Erratum for: On basing one-way functions on NPhardness. In ACM [ACM10], pages 795–796. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x.. See [AGGM06].

Arora:2012:CNM

[AGKM12] Sanjeev Arora, Rong Ge, Ravindran Kannan, and Ankur Moitra. Computing a nonnegative matrix factorization — provably. In ACM [ACM12], pages 145–162. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Andoni:2013:HFU

[AGMP13] Alexandr Andoni, Assaf Goldberger, Andrew McGregor, and Ely Porat. Homomorphic fingerprints under misalignments: sketching edit and shift distances. In ACM [ACM13], pages 931–940. ISBN 1-4503-2029-5.

Adsul:2011:RBG

[AGMS11] Bharat Adsul, Jugal Garg, Ruta Mehta, and Milind Sohoni. Rank-1 bimatrix games: a homeomorphism and a polynomial time algorithm. In ACM [ACM11], pages 195–204. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Ajtai:2010:ORC

[Ajt10] Miklós Ajtai. Oblivious RAMs without cryptographic assumptions. In ACM [ACM10], pages 181–190. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Ajtai:2011:SCI

[Ajt11] Miklos Ajtai. Secure computation with information leaking to an adversary. In ACM [ACM11], pages 715–724. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Ajtai:2012:DVN

[Ajt12] Miklós Ajtai. Determinism versus nondeterminism with arithmetic tests and computation: extended abstract. In ACM [ACM12], pages 249–268. ISBN 1-4503-1245-4. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Ajtai:2013:LBR

[Ajt13] Miklos Ajtai. Lower bounds for RAMs and quantifier elimination. In ACM [ACM13], pages 803–812. ISBN 1-4503-2029-5.

Ambainis:2010:QLL

[AKS10] Andris Ambainis, Julia Kempe, and Or Sattath. A quantum Lovász local lemma. In ACM [ACM10], pages 151–160. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

An:2012:ICA

[AKS12] Hyung-Chan An, Robert Kleinberg, and David B. Shmoys. Improving Christofides’ algorithm for the s-t path TSP. In ACM [ACM12], pages 875–886. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Alon:2013:ARM

[ALSV13] Noga Alon, Troy Lee, Adi Shraibman, and Santosh Vempala. The approximate rank of a matrix and its algorithmic applications: approximate rank. In ACM [ACM13], pages 675–684. ISBN 1-4503-2029-5.

Azar:2012:RP

[AM12] Pablo Daniel Azar and Silvio Micali. Rational proofs. In ACM [ACM12], pages 1017–1028. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Ambainis:2013:SAE

[Amb13] Andris Ambainis. Superlinear advantage for exact quantum algorithms. In ACM [ACM13], pages 891–900. ISBN 1-4503-2029-5.

Alon:2012:NCG

[AMS12] Noga Alon, Ankur Moitra, and Benny Sudakov. Nearly complete graphs decomposable into large induced matchings and their applications. In ACM [ACM12], pages 1079–1090. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Abraham:2012:UPD

[AN12] Ittai Abraham and Ofer Neiman. Using petal-decompositions to build a low stretch spanning tree. In ACM [ACM12], pages 395– 406. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Applebaum:2012:PGL

[App12] Benny Applebaum. Pseudorandom generators with long stretch and low locality from random local one-way functions. In ACM [ACM12], pages 805–816. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Arvind:2010:HND

[AS10] Vikraman Arvind and Srikanth Srinivasan. On the hardness of the noncommutative determinant. In ACM [ACM10], pages 677–686. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Agrawal:2013:QPH

[ASS13] Manindra Agrawal, Chandan Saha, and Nitin Saxena. Quasipolynomial hitting-set for set-depth-Δ formulas. In ACM [ACM13], pages 321–330. ISBN 1-4503-2029-5.

Agrawal:2012:JHC

[ASSS12] Manindra Agrawal, Chandan Saha, Ramprasad Saptharishi, and Nitin Saxena. Jacobian hits circuits: hitting-sets, lower bounds for depth-D occur-k formulas & depth-3 transcendence degree-k circuits. In ACM [ACM12], pages 599–614. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Barak:2010:HCI

[BBCR10] Boaz Barak, Mark Braverman, Xi Chen, and Anup Rao. How to compress interactive communication. In ACM [ACM10], pages 67–76. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Barak:2012:HSS

[BBH^{+}12] Boaz Barak, Fernando G. S. L. Brandao, Aram W. Harrow, Jonathan Kelner, David Steurer, and Yuan Zhou. Hypercontractivity, sum-of-squares proofs, and their applications. In ACM [ACM12], pages 307–326. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Beame:2012:TST

[BBI12] Paul Beame, Christopher Beck, and Russell Impagliazzo. Timespace tradeoffs in resolution: superpolynomial lower bounds for superlinear space. In ACM [ACM12], pages 213–232. ISBN 1-

4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Bansal:2012:TTS

[BBJK12] Nikhil Bansal, Vibhor Bhatt, Prasad Jayanti, and Ranganath Kondapally. Tight time-space tradeoff for mutual exclusion. In ACM [ACM12], pages 971–982. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Burgisser:2010:SPE

[BC10] Peter Bürgisser and Felipe Cucker. Solving polynomial equations in smoothed polynomial time and a near solution to Smale’s 17th problem. In ACM [ACM10], pages 503–512. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Bhaskara:2010:DHL

[BCC^{+}10] Aditya Bhaskara, Moses Charikar, Eden Chlamtac, Uriel Feige, and Aravindan Vijayaraghavan. Detecting high log-densities: an O(n^{1/4}) approximation for densest k-subgraph. In ACM [ACM10], pages 201–210. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010.

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Bitansky:2013:RCB

[BCCT13] Nir Bitansky, Ran Canetti, Alessandro Chiesa, and Eran Tromer. Recursive composition and bootstrapping for SNARKS and proofcarrying data. In ACM [ACM13], pages 111–120. ISBN 1-4503-2029-5.

Bei:2012:BFM

[BCGL12] Xiaohui Bei, Ning Chen, Nick Gravin, and Pinyan Lu. Budget feasible mechanism design: from prior-free to Bayesian. In ACM [ACM12], pages 449–458. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Brandao:2011:QTA

[BCY11] Fernando G. S. L. Brandão, Matthias Christandl, and Jon Yard. A quasipolynomial-time algorithm for the quantum separability problem. In ACM [ACM11], pages 343–352. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Bei:2013:CTE

[BCZ13] Xiaohui Bei, Ning Chen, and Shengyu Zhang. On the complexity of trial and error. In ACM [ACM13], pages 31–40. ISBN 1-4503-2029-5.

Bourgain:2011:BBE

[BDF^{+}11] Jean Bourgain, Stephen J. Dilworth, Kevin Ford, Sergei V. Konyagin, and Denka Kutzarova. Breaking the k^{2} barrier for explicit RIP matrices. In ACM [ACM11], pages 637–644. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Bhaskara:2012:UDP

[BDKT12] Aditya Bhaskara, Daniel Dadush, Ravishankar Krishnaswamy, and Kunal Talwar. Unconditional differentially private mechanisms for linear queries. In ACM [ACM12], pages 1269–1284. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Bhowmick:2013:NBM

[BDL13] Abhishek Bhowmick, Zeev Dvir, and Shachar Lovett. New bounds for matching vector families. In ACM [ACM13], pages 823–832. ISBN 1-4503-2029-5.

Bousquet:2011:MF

[BDT11] Nicolas Bousquet, Jean Daligault, and Stéphan Thomassé. Multicut is FPT. In ACM [ACM11], pages 459–468. ISBN 1-4503-

0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Barak:2011:RBD

[BDYW11] Boaz Barak, Zeev Dvir, Amir Yehudayoff, and Avi Wigderson. Rank bounds for design matrices with applications to combinatorial geometry and locally correctable codes. In ACM [ACM11], pages 519–528. ISBN 1-4503-0691-8. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Belovs:2012:SPF

[Bel12] Aleksandrs Belovs. Span programs for functions with constantsized 1-certificates: extended abstract. In ACM [ACM12], pages 77–84. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Bernstein:2013:MSP

[Ber13] Aaron Bernstein. Maintaining shortest paths under deletions in weighted directed graphs: [extended abstract]. In ACM [ACM13], pages 725–734. ISBN 1-4503-2029-5.

Bhattacharyya:2013:ELC

[BFH^{+}13] Arnab Bhattacharyya, Eldar Fischer, Hamed Hatami, Pooya Hatami, and Shachar Lovett. Every locally characterized affineinvariant property is testable. In ACM [ACM13], pages 429–436. ISBN 1-4503-2029-5.

Bhattacharya:2010:BCA

[BGGM10] Sayan Bhattacharya, Gagan Goel, Sreenivas Gollapudi, and Kamesh Munagala. Budget constrained auctions with heterogeneous items. In ACM [ACM10], pages 379–388. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Bohm:2013:EDO

[BGJ13] Stanislav Böhm, Stefan Göller, and Petr Jancar. Equivalence of deterministic one-counter automata is NL-complete. In ACM [ACM13], pages 131–140. ISBN 1-4503-2029-5.

Boyle:2012:MCS

[BGJK12] Elette Boyle, Shafi Goldwasser, Abhishek Jain, and Yael Tauman Kalai. Multiparty computation secure against continual memory leakage. In ACM [ACM12], pages 1235–1254. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Bartal:2012:TSP

[BGK12] Yair Bartal, Lee-Ad Gottlieb, and Robert Krauthgamer. The traveling salesman problem: low-dimensionality implies a polynomial time approximation scheme. In ACM [ACM12], pages 663–672. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Bhawalkar:2013:COF

[BGM13] Kshipra Bhawalkar, Sreenivas Gollapudi, and Kamesh Munagala. Coevolutionary opinion formation games. In ACM [ACM13], pages 41–50. ISBN 1-4503-2029-5.

Braverman:2013:IEC

[BGPW13] Mark Braverman, Ankit Garg, Denis Pankratov, and Omri Weinstein. From information to exact communication. In ACM [ACM13], pages 151–160. ISBN 1-4503-2029-5.

Byrka:2010:ILB

[BGRS10] Jaroslaw Byrka, Fabrizio Grandoni, Thomas Rothvoß, and Laura Sanità. An improved LP-based approximation for Steiner tree. In ACM [ACM10], pages 583–592. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Bayati:2010:CAI

[BGT10] Mohsen Bayati, David Gamarnik, and Prasad Tetali. Combinatorial approach to the interpolation method and scaling limits in sparse random graphs. In ACM [ACM10], pages 105–114. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Balcan:2011:LSF

[BH11] Maria-Florina Balcan and Nicholas J. A. Harvey. Learning submodular functions. In ACM [ACM11], pages 793–802. ISBN 1-

4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Brandao:2013:PSA

[BH13a] Fernando G. S. L. Brandao and Aram W. Harrow. Product-state approximations to quantum ground states. In ACM [ACM13], pages 871–880. ISBN 1-4503-2029-5.

Brandao:2013:QFT

[BH13b] Fernando G. S. L. Brandao and Aram W. Harrow. Quantum de Finetti theorems under local measurements with applications. In ACM [ACM13], pages 861–870. ISBN 1-4503-2029-5.

Bateni:2010:ASS

[BHM10] MohammadHossein Bateni, MohammadTaghi Hajiaghayi, and Dániel Marx. Approximation schemes for Steiner forest on planar graphs and graphs of bounded treewidth. In ACM [ACM10], pages 211–220. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010.

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Beame:2010:HAP

[BHP10] Paul Beame, Trinh Huynh, and Toniann Pitassi. Hardness amplification in proof complexity. In ACM [ACM10], pages 87–96. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Burgisser:2011:GCT

[BI11] Peter Bürgisser and Christian Ikenmeyer. Geometric complexity theory and tensor rank. In ACM [ACM11], pages 509–518. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Beck:2013:SEH

[BI13a] Christopher Beck and Russell Impagliazzo. Strong ETH holds for regular resolution. In ACM [ACM13], pages 487–494. ISBN 1-4503-2029-5.

Burgisser:2013:ELB

[BI13b] Peter Bürgisser and Christian Ikenmeyer. Explicit lower bounds via geometric complexity theory. In ACM [ACM13], pages 141– 150. ISBN 1-4503-2029-5.

Brandt:2012:AOD

[BIKK12] Christina Brandt, Nicole Immorlica, Gautam Kamath, and Robert Kleinberg. An analysis of one-dimensional Schelling segregation. In ACM [ACM12], pages 789–804. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Barto:2012:RSC

[BK12] Libor Barto and Marcin Kozik. Robust satisfiability of constraint satisfaction problems. In ACM [ACM12], pages 931–940. ISBN

1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Bulnek:2012:TLB

[BKS12] Jan Bulánek, Michal Koucký, and Michael Saks. Tight lower bounds for the online labeling problem. In ACM [ACM12], pages 1185–1198. ISBN 1-4503-1245-4. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Bringmann:2013:SSD

[BL13] Karl Bringmann and Kasper Green Larsen. Succinct sampling from discrete distributions. In ACM [ACM13], pages 775–782. ISBN 1-4503-2029-5.

Brakerski:2013:CHL

[BLP^{+}13] Zvika Brakerski, Adeline Langlois, Chris Peikert, Oded Regev, and Damien Stehlé. Classical hardness of learning with errors. In ACM [ACM13], pages 575–584. ISBN 1-4503-2029-5.

Brodal:2012:SFH

[BLT12] Gerth Stølting Brodal, George Lagogiannis, and Robert E. Tarjan. Strict Fibonacci heaps. In ACM [ACM12], pages 1177–1184. ISBN

1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Braverman:2013:ICA

[BM13] Mark Braverman and Ankur Moitra. An information complexity approach to extended formulations. In ACM [ACM13], pages 161– 170. ISBN 1-4503-2029-5.

Buchbinder:2013:SPE

[BNS13] Niv Buchbinder, Joseph (Seffi) Naor, and Roy Schwartz. Simplex partitioning via exponential clocks and the multiway cut problem. In ACM [ACM13], pages 535–544. ISBN 1-4503-2029-5.

Beck:2013:STR

[BNT13] Chris Beck, Jakob Nordstrom, and Bangsheng Tang. Some tradeoff results for polynomial calculus: extended abstract. In ACM [ACM13], pages 813–822. ISBN 1-4503-2029-5.

Braverman:2010:MID

[BO10a] Vladimir Braverman and Rafail Ostrovsky. Measuring independence of datasets. In ACM [ACM10], pages 271–280. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Braverman:2010:ZOF

[BO10b] Vladimir Braverman and Rafail Ostrovsky. Zero-one frequency laws. In ACM [ACM10], pages 281–290. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Bodirsky:2011:STG

[BP11] Manuel Bodirsky and Michael Pinsker. Schaefer’s theorem for graphs. In ACM [ACM11], pages 655–664. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Bitansky:2013:IAO

[BP13] Nir Bitansky and Omer Paneth. On the impossibility of approximate obfuscation and applications to resettable cryptography. In ACM [ACM13], pages 241–250. ISBN 1-4503-2029-5.

Braverman:2011:TCM

[BR11] Mark Braverman and Anup Rao. Towards coding for maximum errors in interactive communication. In ACM [ACM11], pages 159–166. ISBN 1-4503-0691-8. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Brunsch:2012:ISA

[BR12] Tobias Brunsch and Heiko Röglin. Improved smoothed analysis of multiobjective optimization. In ACM [ACM12], pages 407–426. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Braverman:2012:IIC

[Bra12] Mark Braverman. Interactive information complexity. In ACM [ACM12], pages 505–524. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Ben-Sasson:2013:CEP

[BSCGT13] Eli Ben-Sasson, Alessandro Chiesa, Daniel Genkin, and Eran Tromer. On the concrete efficiency of probabilistically-checkable proofs. In ACM [ACM13], pages 585–594. ISBN 1-4503-2029-5.

Ben-Sasson:2013:NFL

[BSGK^{+}13] Eli Ben-Sasson, Ariel Gabizon, Yohay Kaplan, Swastik Kopparty, and Shubangi Saraf. A new family of locally correctable codes based on degree-lifted algebraic geometry codes. In ACM [ACM13], pages 833–842. ISBN 1-4503-2029-5.

Babai:2013:QTC

[BW13] Laszlo Babai and John Wilmes. Quasipolynomial-time canonical form for Steiner designs. In ACM [ACM13], pages 261–270. ISBN 1-4503-2029-5.

Bourgain:2012:ME

[BY12] Jean Bourgain and Amir Yehudayoff. Monotone expansion. In ACM [ACM12], pages 1061–1078. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Cai:2012:CCC

[CC12] Jin-Yi Cai and Xi Chen. Complexity of counting CSP with complex weights. In ACM [ACM12], pages 909–920. ISBN 1-4503-

1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Chekuri:2013:LTG

[CC13] Chandra Chekuri and Julia Chuzhoy. Large-treewidth graph decompositions and applications. In ACM [ACM13], pages 291–300. ISBN 1-4503-2029-5.

Cheung:2013:TBG

[CCD13] Yun Kuen Cheung, Richard Cole, and Nikhil Devanur. Tatonnement beyond gross substitutes?: gradient descent to the rescue. In ACM [ACM13], pages 191–200. ISBN 1-4503-2029-5.

Chalermsook:2012:AAH

[CCEL12] Parinya Chalermsook, Julia Chuzhoy, Alina Ene, and Shi Li. Approximation algorithms and hardness of integral concurrent flow. In ACM [ACM12], pages 689–708. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Cole:2011:IPS

[CCG^{+}11] Richard Cole, José R. Correa, Vasilis Gkatzelis, Vahab Mirrokni, and Neil Olver. Inner product spaces for MinSum coordination mechanisms. In ACM [ACM11], pages 539–548. ISBN 1-4503-

0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Cebrian:2012:FRB

[CCVV12] Manuel Cebrian, Lorenzo Coviello, Andrea Vattani, and Panagiotis Voulgaris. Finding red balloons with split contracts: robustness to individuals’ selfishness. In ACM [ACM12], pages 775–788. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Cai:2012:ACM

[CDW12] Yang Cai, Constantinos Daskalakis, and S. Matthew Weinberg. An algorithmic characterization of multi-dimensional mechanisms. In ACM [ACM12], pages 459–478. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Cardinal:2010:SUP

[CFJ^{+}10] Jean Cardinal, Samuel Fiorini, Gwenaël Joret, Raphaël M. Jungers, and J. Ian Munro. Sorting under partial information (without the ellipsoid algorithm). In ACM [ACM10], pages 359– 368. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Chakrabarti:2012:WCC

[CFW12] Amit Chakrabarti, Lisa Fleischer, and Christophe Weibel. When the cut condition is enough: a complete characterization for multiflow problems in series-parallel networks. In ACM [ACM12], pages 19–26. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Cai:2013:CDR

[CGW13] Jin-Yi Cai, Heng Guo, and Tyson Williams. A complete dichotomy rises from the capture of vanishing signatures: extended abstract. In ACM [ACM13], pages 635–644. ISBN 1-4503-2029-5.

Chan:2013:ARP

[Cha13] Siu On Chan. Approximation resistance from pairwise independent subgroups. In ACM [ACM13], pages 447–456. ISBN 1-4503-2029-5.

Censor-Hillel:2012:GCP

[CHHKM12] Keren Censor-Hillel, Bernhard Haeupler, Jonathan Kelner, and Petar Maymounkov. Global computation in a poorly connected world: fast rumor spreading with no dependence on conductance. In ACM [ACM12], pages 961–970. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Chawla:2010:MPM

[CHMS10] Shuchi Chawla, Jason D. Hartline, David L. Malec, and Balasubramanian Sivan. Multi-parameter mechanism design and sequential posted pricing. In ACM [ACM10], pages 311–320. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Chawla:2013:PIM

[CHMS13] Shuchi Chawla, Jason D. Hartline, David Malec, and Balasubramanian Sivan. Prior-independent mechanisms for scheduling. In ACM [ACM13], pages 51–60. ISBN 1-4503-2029-5.

Chien:2011:ASH

[CHSS11] Steve Chien, Prahladh Harsha, Alistair Sinclair, and Srikanth Srinivasan. Almost settling the hardness of noncommutative determinant. In ACM [ACM11], pages 499–508. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Chuzhoy:2011:AGC

[Chu11] Julia Chuzhoy. An algorithm for the graph crossing number problem. In ACM [ACM11], pages 303–312. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Chuzhoy:2012:VSS

[Chu12a] Julia Chuzhoy. On vertex sparsifiers with Steiner nodes. In ACM [ACM12], pages 673–688. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Chuzhoy:2012:RUG

[Chu12b] Julia Chuzhoy. Routing in undirected graphs with constant congestion. In ACM [ACM12], pages 855–874. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Chawla:2012:LBB

[CIL12] Shuchi Chawla, Nicole Immorlica, and Brendan Lucier. On the limits of black-box reductions in mechanism design. In ACM [ACM12], pages 435–448. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Cheung:2012:FMR

[CKL12] Ho Yee Cheung, Tsz Chiu Kwok, and Lap Chi Lau. Fast matrix rank algorithms and applications. In ACM [ACM12], pages 549– 562. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Christiano:2011:EFL

[CKM^{+}11] Paul Christiano, Jonathan A. Kelner, Aleksander Madry, Daniel A. Spielman, and Shang-Hua Teng. Electrical flows, Laplacian systems, and faster approximation of maximum flow in undirected graphs. In ACM [ACM11], pages 273–282. ISBN 1-4503-

0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Cadek:2013:ECM

[CKM^{+}13] Martin Cadek, Marek Krcal, Jiri Matousek, Lukas Vokrinek, and Uli Wagner. Extending continuous maps: polynomiality and undecidability. In ACM [ACM13], pages 595–604. ISBN 1-4503-2029-5.

Cygan:2013:FHC

[CKN13] Marek Cygan, Stefan Kratsch, and Jesper Nederlof. Fast Hamiltonicity checking via bases of perfect matchings. In ACM [ACM13], pages 301–310. ISBN 1-4503-2029-5.

Chandran:2010:PAA

[CKOR10] Nishanth Chandran, Bhavana Kanukurthi, Rafail Ostrovsky, and Leonid Reyzin. Privacy amplification with asymptotically optimal entropy loss. In ACM [ACM10], pages 785–794. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Chierichetti:2010:ATB

[CLP10] Flavio Chierichetti, Silvio Lattanzi, and Alessandro Panconesi. Almost tight bounds for rumour spreading with conductance. In ACM [ACM10], pages 399–408. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Caputo:2013:RLT

[CMSS13] Pietro Caputo, Fabio Martinelli, Alistair Sinclair, and Alexandre Stauffer. Random lattice triangulations: structure and algorithms. In ACM [ACM13], pages 615–624. ISBN 1-4503-2029-5.

Coja-Oglan:2012:CKN

[COP12] Amin Coja-Oglan and Konstantinos Panagiotou. Catching the k-NAESAT threshold. In ACM [ACM12], pages 899–908. ISBN

1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Coja-Oghlan:2013:GAK

[COP13] Amin Coja-Oghlan and Konstantinos Panagiotou. Going after the k-SAT threshold. In ACM [ACM13], pages 705–714. ISBN 1-4503-2029-5.

Chonev:2013:OPH

[COW13] Ventsislav Chonev, Joël Ouaknine, and James Worrell. The orbit problem in higher dimensions. In ACM [ACM13], pages 941–950. ISBN 1-4503-2029-5.

Chan:2012:TBM

[CP12] Siu Man Chan and Aaron Potechin. Tight bounds for monotone switching networks via Fourier analysis. In ACM [ACM12], pages 495–504. ISBN 1-4503-1245-4. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Chung:2013:NBB

[CPS13] Kai-Min Chung, Rafael Pass, and Karn Seth. Non-black-box simulation from one-way functions and applications to resettable security. In ACM [ACM13], pages 231–240. ISBN 1-4503-2029-5.

Chen:2013:CNM

[CPY13] Xi Chen, Dimitris Paparas, and Mihalis Yannakakis. The complexity of non-monotone markets. In ACM [ACM13], pages 181–190. ISBN 1-4503-2029-5.

Chakrabarti:2011:OLB

[CR11] Amit Chakrabarti and Oded Regev. An optimal lower bound on the communication complexity of gap-Hamming-distance. In ACM [ACM11], pages 51–60. ISBN 1-4503-0691-8. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Chakrabarty:2013:MTB

[CS13a] Deeparnab Chakrabarty and C. Seshadhri. A o(n) monotonicity tester for Boolean functions over the hypercube. In ACM [ACM13], pages 411–418. ISBN 1-4503-2029-5.

Chakrabarty:2013:OBM

[CS13b] Deeparnab Chakrabarty and C. Seshadhri. Optimal bounds for monotonicity and Lipschitz testing over hypercubes and hypergrids. In ACM [ACM13], pages 419–428. ISBN 1-4503-2029-5.

Chen:2013:MSD

[CST13] Xi Chen, Xiaorui Sun, and Shang-Hua Teng. Multi-stage design for quasipolynomial-time isomorphism testing of Steiner 2systems. In ACM [ACM13], pages 271–280. ISBN 1-4503-2029-5.

Clarkson:2013:LRA

[CW13] Kenneth L. Clarkson and David P. Woodruff. Low rank approximation and regression in input sparsity time. In ACM [ACM13], pages 81–90. ISBN 1-4503-2029-5.

De:2012:NOS

[DDFS12] Anindya De, Ilias Diakonikolas, Vitaly Feldman, and Rocco A. Servedio. Nearly optimal solutions for the chow parameters problem and low-weight approximation of halfspaces. In ACM [ACM12], pages 729–746. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Daskalakis:2012:LPB

[DDS12] Constantinos Daskalakis, Ilias Diakonikolas, and Rocco A. Servedio. Learning Poisson binomial distributions. In ACM [ACM12], pages 709–728. ISBN 1-4503-1245-4. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Doerr:2011:SNS

[DFF11] Benjamin Doerr, Mahmoud Fouz, and Tobias Friedrich. Social networks spread rumors in sublogarithmic time. In ACM [ACM11], pages 21–30. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Dobzinski:2011:OAC

[DFK11] Shahar Dobzinski, Hu Fu, and Robert D. Kleinberg. Optimal auctions with correlated bidders are easy. In ACM [ACM11], pages 129–138. ISBN 1-4503-0691-8. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Demaine:2011:CDH

[DHiK11] Erik D. Demaine, MohammadTaghi Hajiaghayi, and Ken ichi Kawarabayashi. Contraction decomposition in h-minor-free graphs and algorithmic applications. In ACM [ACM11], pages 441–450. ISBN 1-4503-0691-8. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Dey:2010:OHC

[DHK10a] Tamal K. Dey, Anil N. Hirani, and Bala Krishnamoorthy. Optimal homologous cycles, total unimodularity, and linear programming. In ACM [ACM10], pages 221–230. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Diakonikolas:2010:BAS

[DHK^{+}10b] Ilias Diakonikolas, Prahladh Harsha, Adam Klivans, Raghu Meka, Prasad Raghavendra, Rocco A. Servedio, and Li-Yang Tan. Bounding the average sensitivity and noise sensitivity of polynomial threshold functions. In ACM [ACM10], pages 533–542. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Devanur:2012:OMC

[DJ12] Nikhil R. Devanur and Kamal Jain. Online matching with concave returns. In ACM [ACM12], pages 137–144. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Dinitz:2011:DSF

[DK11] Michael Dinitz and Robert Krauthgamer. Directed spanners via flow-based linear programs. In ACM [ACM11], pages 323–332. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Dasgupta:2010:SJL

[DKS10] Anirban Dasgupta, Ravi Kumar, and Tamás Sarlos. A sparse Johnson–Lindenstrauss transform. In ACM [ACM10], pages 341– 350. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Dvir:2012:SES

[DL12] Zeev Dvir and Shachar Lovett. Subspace evasive sets. In ACM [ACM12], pages 351–358. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Ding:2011:CTB

[DLP11] Jian Ding, James R. Lee, and Yuval Peres. Cover times, blanket times, and majorizing measures. In ACM [ACM11], pages 61–70. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

De:2013:MSD

[DMN13] Anindya De, Elchanan Mossel, and Joe Neeman. Majority is stablest: discrete and SoS. In ACM [ACM13], pages 477–486. ISBN 1-4503-2029-5.

Dvir:2012:SMB

[DMPY12] Zeev Dvir, Guillaume Malod, Sylvain Perifel, and Amir Yehudayoff. Separating multilinear branching programs and formulas. In ACM [ACM12], pages 615–624. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Dwork:2010:DPU

[DNPR10] Cynthia Dwork, Moni Naor, Toniann Pitassi, and Guy N. Rothblum. Differential privacy under continual observation. In ACM [ACM10], pages 715–724. ISBN 1-60558-817-2. LCCN QA 76.6

.A152 2010. URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Dobzinski:2011:IRT

[Dob11] Shahar Dobzinski. An impossibility result for truthful combinatorial auctions with submodular valuations. In ACM [ACM11], pages 139–148. ISBN 1-4503-0691-8. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Duan:2010:COF

[DP10] Ran Duan and Seth Pettie. Connectivity oracles for failure prone graphs. In ACM [ACM10], pages 465–474. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Dodis:2010:CBL

[DPT10] Yevgeniy Dodis, Mihai Patrascu, and Mikkel Thorup. Changing base without losing space. In ACM [ACM10], pages 593–602. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Dyer:2010:CC

[DR10] Martin E. Dyer and David M. Richerby. On the complexity of #csp. In ACM [ACM10], pages 725–734. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Dughmi:2011:COR

[DRY11] Shaddin Dughmi, Tim Roughgarden, and Qiqi Yan. From convex optimization to randomized mechanisms: toward optimal combinatorial auctions. In ACM [ACM11], pages 149–158. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

De:2010:NOE

[DV10] Anindya De and Thomas Vidick. Near-optimal extractors against quantum storage. In ACM [ACM10], pages 161–170. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Dobzinski:2012:QCC

[DV12] Shahar Dobzinski and Jan Vondrak. From query complexity to computational complexity. In ACM [ACM12], pages 1107–1116.

ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Dell:2010:SAN

[DvM10] Holger Dell and Dieter van Melkebeek. Satisfiability allows no nontrivial sparsification unless the polynomial-time hierarchy collapses. In ACM [ACM10], pages 251–260. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Efremenko:2012:IRL

[Efr12] Klim Efremenko. From irreducible representations to locally decodable codes. In ACM [ACM12], pages 327–338. ISBN 1-4503-

1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Eisenstat:2013:LTA

[EK13] David Eisenstat and Philip N. Klein. Linear-time algorithms for max flow and multiple-source shortest paths in unit-weight planar graphs. In ACM [ACM13], pages 735–744. ISBN 1-4503-2029-5.

Elkin:2013:OES

[ES13] Michael Elkin and Shay Solomon. Optimal Euclidean spanners: really short, thin and lanky. In ACM [ACM13], pages 645–654. ISBN 1-4503-2029-5.

Etessami:2012:PTA

[ESY12] Kousha Etessami, Alistair Stewart, and Mihalis Yannakakis. Polynomial time algorithms for multi-type branching processes and stochastic context-free grammars. In ACM [ACM12], pages 579– 588. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Feldman:2013:SAA

[FFGL13] Michal Feldman, Hu Fu, Nick Gravin, and Brendan Lucier. Simultaneous auctions are (almost) efficient. In ACM [ACM13], pages 201–210. ISBN 1-4503-2029-5.

Fraigniaud:2010:SSW

[FG10] Pierre Fraigniaud and George Giakkoupis. On the searchability of small-world networks with arbitrary underlying structure. In ACM [ACM10], pages 389–398. ISBN 1-60558-817-2. LCCN QA 76.6

.A152 2010. URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Feldman:2013:CWE

[FGL13] Michal Feldman, Nick Gravin, and Brendan Lucier. Combinatorial Walrasian Equilibrium. In ACM [ACM13], pages 61–70. ISBN 1-4503-2029-5.

Feldman:2013:SAL

[FGR^{+}13] Vitaly Feldman, Elena Grigorescu, Lev Reyzin, Santosh Vempala, and Ying Xiao. Statistical algorithms and a lower bound for detecting planted cliques. In ACM [ACM13], pages 655–664. ISBN 1-4503-2029-5.

Fung:2011:GFG

[FHHP11] Wai Shing Fung, Ramesh Hariharan, Nicholas J. A. Harvey, and Debmalya Panigrahi. A general framework for graph sparsification. In ACM [ACM11], pages 71–80. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Fawzi:2011:LDN

[FHS11] Omar Fawzi, Patrick Hayden, and Pranab Sen. From lowdistortion norm embeddings to explicit uncertainty relations and efficient information locking. In ACM [ACM11], pages 773–782. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Friedmann:2011:SLB

[FHZ11] Oliver Friedmann, Thomas Dueholm Hansen, and Uri Zwick. Subexponential lower bounds for randomized pivoting rules for the simplex algorithm. In ACM [ACM11], pages 283–292. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Fraigniaud:2010:OAS

[FK10] Pierre Fraigniaud and Amos Korman. An optimal ancestry scheme and small universal posets. In ACM [ACM10], pages 611–620. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Feldman:2011:UFA

[FL11] Dan Feldman and Michael Langberg. A unified framework for approximating and clustering data. In ACM [ACM11], pages 569– 578. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Fiorini:2012:LVS

[FMP^{+}12] Samuel Fiorini, Serge Massar, Sebastian Pokutta, Hans Raj Tiwary, and Ronald de Wolf. Linear vs. semidefinite extended formulations: exponential separation and strong lower bounds. In ACM [ACM12], pages 95–106. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Forbes:2012:ITT

[FS12] Michael A. Forbes and Amir Shpilka. On identity testing of tensors, low-rank recovery and compressed sensing. In ACM [ACM12], pages 163–172. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Feige:2011:MDU

[FT11] Uriel Feige and Moshe Tennenholtz. Mechanism design with uncertain inputs: (to err is human, to forgive divine). In ACM [ACM11], pages 549–558. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Gu:2013:PDM

[GGK13] Albert Gu, Anupam Gupta, and Amit Kumar. The power of deferral: maintaining a constant-competitive Steiner tree online. In ACM [ACM13], pages 525–534. ISBN 1-4503-2029-5.

Godoy:2010:HPD

[GGRÀ10] Guillem Godoy, Omer Giménez, Lander Ramos, and Carme Àlvarez. The HOM problem is decidable. In ACM [ACM10], pages 485–494. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010.

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Garg:2013:WEA

[GGSW13] Sanjam Garg, Craig Gentry, Amit Sahai, and Brent Waters. Witness encryption and its applications. In ACM [ACM13], pages 467–476. ISBN 1-4503-2029-5.

Guruswami:2010:LDR

[GHK10] Venkatesan Guruswami, Johan Håstad, and Swastik Kopparty. On the list-decodability of random linear codes. In ACM [ACM10], pages 409–416. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010.

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Gal:2012:TBC

[GHK^{+}12] Anna Gál, Kristoffer Arnsfelt Hansen, Michal Koucký, Pavel Pudlák, and Emanuele Viola. Tight bounds on computing error-correcting codes by bounded-depth circuits with arbitrary gates. In ACM [ACM12], pages 479–494. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Gupta:2011:PRC

[GHRU11] Anupam Gupta, Moritz Hardt, Aaron Roth, and Jonathan Ullman. Privately releasing conjunctions and the statistical query barrier. In ACM [ACM11], pages 803–812. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Golab:2011:LID

[GHW11] Wojciech Golab, Lisa Higham, and Philipp Woelfel. Linearizable implementations do not suffice for randomized distributed computation. In ACM [ACM11], pages 373–382. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Grohe:2011:FTS

[GiKMW11] Martin Grohe, Ken ichi Kawarabayashi, Dániel Marx, and Paul Wollan. Finding topological subgraphs is fixed-parameter tractable. In ACM [ACM11], pages 479–488. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Goyal:2010:RCC

[GJ10] Vipul Goyal and Abhishek Jain. On the round complexity of covert computation. In ACM [ACM10], pages 191–200. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Goyal:2012:CCN

[GK12] Sanjeev Goyal and Michael Kearns. Competitive contagion in networks. In ACM [ACM12], pages 759–774. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Goel:2010:PML

[GKK10] Ashish Goel, Michael Kapralov, and Sanjeev Khanna. Perfect matchings in o(n log n) time in regular bipartite graphs. In ACM [ACM10], pages 39–46. ISBN 1-60558-817-2. LCCN QA 76.6

.A152 2010. URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Gupta:2012:RDM

[GKL12] Ankit Gupta, Neeraj Kayal, and Satya Lokam. Reconstruction of depth-4 multilinear circuits with top fan-in 2. In ACM [ACM12], pages 625–642. ISBN 1-4503-1245-4. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Goldwasser:2013:RGC

[GKP^{+}13] Shafi Goldwasser, Yael Kalai, Raluca Ada Popa, Vinod Vaikuntanathan, and Nickolai Zeldovich. Reusable garbled circuits and succinct functional encryption. In ACM [ACM13], pages 555–564. ISBN 1-4503-2029-5.

Gilbert:2010:ASR

[GLPS10] Anna C. Gilbert, Yi Li, Ely Porat, and Martin J. Strauss. Approximate sparse recovery: optimizing time and measurements. In ACM [ACM10], pages 475–484. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Grohe:2012:STI

[GM12] Martin Grohe and Dániel Marx. Structure theorem and isomorphism test for graphs with excluded topological subgraphs. In ACM [ACM12], pages 173–192. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Goel:2012:PCA

[GML12] Gagan Goel, Vahab Mirrokni, and Renato Paes Leme. Polyhedral clinching auctions and the AdWords polytope. In ACM [ACM12], pages 107–122. ISBN 1-4503-1245-4. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Gopalan:2011:PGC

[GMRZ11] Parikshit Gopalan, Raghu Meka, Omer Reingold, and David Zuckerman. Pseudorandom generators for combinatorial shapes. In ACM [ACM11], pages 253–262. ISBN 1-4503-0691-8. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Garg:2012:CPA

[GMSV12] Jugal Garg, Ruta Mehta, Milind Sohoni, and Vijay V. Vazirani. A complementary pivot algorithm for markets under separable, piecewise-linear concave utilities. In ACM [ACM12], pages 1003– 1016. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Goemans:2012:MIG

[GORZ12] Michel X. Goemans, Neil Olver, Thomas Rothvoß, and Rico Zenklusen. Matroids and integrality gaps for hypergraphic Steiner tree relaxations. In ACM [ACM12], pages 1161–1176. ISBN 1-

4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Goyal:2011:CRN

[Goy11] Vipul Goyal. Constant round non-malleable protocols using one way functions. In ACM [ACM11], pages 695–704. ISBN 1-4503-

0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Goyal:2013:NBB

[Goy13] Vipul Goyal. Non-black-box simulation in the fully concurrent setting. In ACM [ACM13], pages 221–230. ISBN 1-4503-2029-5.

Giakkoupis:2011:OPS

[GS11] George Giakkoupis and Nicolas Schabanel. Optimal path search in small worlds: dimension matters. In ACM [ACM11], pages 393–402. ISBN 1-4503-0691-8. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Gupta:2013:SCB

[GTW13] Anupam Gupta, Kunal Talwar, and David Witmer. Sparsest cut on bounded treewidth graphs: algorithms and hardness results. In ACM [ACM13], pages 281–290. ISBN 1-4503-2029-5.

Gorbunov:2013:ABE

[GVW13] Sergey Gorbunov, Vinod Vaikuntanathan, and Hoeteck Wee. Attribute-based encryption for circuits. In ACM [ACM13], pages 545–554. ISBN 1-4503-2029-5.

Gao:2010:LBO

[GW10] Pu Gao and Nicholas C. Wormald. Load balancing and orientability thresholds for random hypergraphs. In ACM [ACM10], pages 97–104. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Gentry:2011:SSN

[GW11] Craig Gentry and Daniel Wichs. Separating succinct noninteractive arguments from all falsifiable assumptions. In ACM [ACM11], pages 99–108. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Giakkoupis:2012:TRL

[GW12] George Giakkoupis and Philipp Woelfel. A tight RMR lower bound for randomized mutual exclusion. In ACM [ACM12], pages 983– 1002. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Guruswami:2012:FCF

[GX12] Venkatesan Guruswami and Chaoping Xing. Folded codes from function field towers and improved optimal rate list decoding. In ACM [ACM12], pages 339–350. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Guruswami:2013:LDR

[GX13] Venkatesan Guruswami and Chaoping Xing. List decoding Reed– Solomon, algebraic-geometric, and Gabidulin subcodes up to the singleton bound. In ACM [ACM13], pages 843–852. ISBN 1-4503-2029-5.

Haeupler:2011:ANC

[Hae11] Bernhard Haeupler. Analyzing network coding gossip made easy. In ACM [ACM11], pages 293–302. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Hassanieh:2012:NOS

[HIKP12] Haitham Hassanieh, Piotr Indyk, Dina Katabi, and Eric Price. Nearly optimal sparse Fourier transform. In ACM [ACM12], pages 563–578. ISBN 1-4503-1245-4. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Hirai:2010:MMP

[Hir10] Hiroshi Hirai. The maximum multiflow problems with bounded fractionality. In ACM [ACM10], pages 115–120. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Heilman:2012:SPC

[HJN12] Steven Heilman, Aukosh Jagannath, and Assaf Naor. Solution of the propeller conjecture in R^{3}. In ACM [ACM12], pages 269–276. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Hansen:2011:EAS

[HKL^{+}11] Kristoffer Arnsfelt Hansen, Michal Koucky, Niels Lauritzen, Peter Bro Miltersen, and Elias P. Tsigaridas. Exact algorithms for solving stochastic games: extended abstract. In ACM [ACM11], pages 205–214. ISBN 1-4503-0691-8. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Harsha:2010:IPP

[HKM10] Prahladh Harsha, Adam Klivans, and Raghu Meka. An invariance principle for polytopes. In ACM [ACM10], pages 543–552. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Holenstein:2011:ERO

[HKT11] Thomas Holenstein, Robin Künzler, and Stefano Tessaro. The equivalence of the random oracle model and the ideal cipher model, revisited. In ACM [ACM11], pages 89–98. ISBN 1-4503-

0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Hartline:2010:BAM

[HL10] Jason D. Hartline and Brendan Lucier. Bayesian algorithmic mechanism design. In ACM [ACM10], pages 301–310. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Hatami:2011:CTA

[HL11] Hamed Hatami and Shachar Lovett. Correlation testing for affine invariant properties on F_{p}^{n} in the high error regime. In ACM [ACM11], pages 187–194. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Huynh:2012:VSP

[HN12] Trinh Huynh and Jakob Nordstrom. On the virtue of succinct proofs: amplifying communication complexity hardness to timespace trade-offs in proof complexity. In ACM [ACM12], pages 233–248. ISBN 1-4503-1245-4. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Har-Peled:2013:NPL

[HPR13] Sariel Har-Peled and Benjamin Adam Raichel. Net and prune: a linear time algorithm for Euclidean distance problems. In ACM [ACM13], pages 605–614. ISBN 1-4503-2029-5.

Hardt:2012:BRR

[HR12] Moritz Hardt and Aaron Roth. Beating randomized response on incoherent matrices. In ACM [ACM12], pages 1255–1268. ISBN

1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Hardt:2013:BWC

[HR13] Moritz Hardt and Aaron Roth. Beyond worst-case analysis in private singular vector computation. In ACM [ACM13], pages 331–340. ISBN 1-4503-2029-5.

Hsu:2013:DPA

[HRU13] Justin Hsu, Aaron Roth, and Jonathan Ullman. Differential privacy for the analyst via private equilibrium computation. In ACM [ACM13], pages 341–350. ISBN 1-4503-2029-5.

Haitner:2010:EIC

[HRV10] Iftach Haitner, Omer Reingold, and Salil Vadhan. Efficiency improvements in constructing pseudorandom generators from oneway functions. In ACM [ACM10], pages 437–446. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Haramaty:2010:SCQ

[HS10] Elad Haramaty and Amir Shpilka. On the structure of cubic and quartic polynomials. In ACM [ACM10], pages 331–340. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Harris:2013:CSP

[HS13] David G. Harris and Aravind Srinivasan. Constraint satisfaction, packet routing, and the Lovasz local lemma. In ACM [ACM13], pages 685–694. ISBN 1-4503-2029-5.

Hardt:2010:GDP

[HT10] Moritz Hardt and Kunal Talwar. On the geometry of differential privacy. In ACM [ACM10], pages 705–714. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Hrubes:2012:SPD

[HT12] Pavel Hrubes and Iddo Tzameret. Short proofs for the determinant identities. In ACM [ACM12], pages 193–212. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Huang:2013:ARS

[Hua13] Sangxia Huang. Approximation resistance on satisfiable instances for predicates with few accepting inputs. In ACM [ACM13], pages 457–466. ISBN 1-4503-2029-5.

Hardt:2013:HRL

[HW13] Moritz Hardt and David P. Woodruff. How robust are linear sketches to adaptive inputs? In ACM [ACM13], pages 121–130. ISBN 1-4503-2029-5.

HrubeAa:2010:NCC

[HWY10] Pavel Hrube, Avi Wigderson, and Amir Yehudayoff. Noncommutative circuits and the sum-of-squares problem. In ACM [ACM10], pages 667–676. ISBN 1-60558-817-2. LCCN QA 76.6

.A152 2010. URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Kawarabayashi:2011:BAA

[iKK11] Ken ichi Kawarabayashi and Yusuke Kobayashi. Breaking o(n^{1/2})- approximation algorithms for the edge-disjoint paths problem with congestion two. In ACM [ACM11], pages 81–88. ISBN 1-4503-

0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Immorlica:2011:DA

[IKL^{+}11] Nicole Immorlica, Adam Tauman Kalai, Brendan Lucier, Ankur Moitra, Andrew Postlewaite, and Moshe Tennenholtz. Dueling algorithms. In ACM [ACM11], pages 215–224. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Kawarabayashi:2010:OCP

[iKR10] Ken ichi Kawarabayashi and Bruce Reed. Odd cycle packing. In ACM [ACM10], pages 695–704. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Kawarabayashi:2010:SPG

[iKW10] Ken ichi Kawarabayashi and Paul Wollan. A shorter proof of the graph minor algorithm: the unique linkage theorem. In ACM [ACM10], pages 687–694. ISBN 1-60558-817-2. LCCN QA 76.6

.A152 2010. URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Kawarabayashi:2011:SAS

[iKW11] Ken ichi Kawarabayashi and Paul Wollan. A simpler algorithm and shorter proof for the graph minor decomposition. In ACM [ACM11], pages 451–458. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Kawarabayashi:2013:TSF

[iKY13] Ken ichi Kawarabayashi and Yuichi Yoshida. Testing subdivisionfreeness: property testing meets structural graph theory. In ACM [ACM13], pages 437–446. ISBN 1-4503-2029-5.

Italiano:2011:IAM

[INSWN11] Giuseppe F. Italiano, Yahav Nussbaum, Piotr Sankowski, and Christian Wulff-Nilsen. Improved algorithms for min cut and max flow in undirected planar graphs. In ACM [ACM11], pages 313– 322. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Indyk:2011:KMC

[IP11] Piotr Indyk and Eric Price. K-median clustering, model-based compressive sensing, and sparse recovery for earth mover distance. In ACM [ACM11], pages 627–636. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Jain:2010:QP

[JJUW10] Rahul Jain, Zhengfeng Ji, Sarvagya Upadhyay, and John Watrous. QIP = PSPACE. In ACM [ACM10], pages 573–582. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Jain:2013:LRM

[JNS13] Prateek Jain, Praneeth Netrapalli, and Sujay Sanghavi. Lowrank matrix completion using alternating minimization. In ACM [ACM13], pages 665–674. ISBN 1-4503-2029-5.

Kannan:2010:SMM

[Kan10] Ravindran Kannan. Spectral methods for matrices and tensors. In ACM [ACM10], pages 1–12. ISBN 1-60558-817-2. LCCN QA 76.6

.A152 2010. URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Karnin:2011:DCH

[Kar11] Zohar S. Karnin. Deterministic construction of a high dimensional l_{p} section in l_{1}^{n} for any p < 2. In ACM [ACM11], pages 645–654. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Kayal:2012:APP

[Kay12] Neeraj Kayal. Affine projections of polynomials: extended abstract. In ACM [ACM12], pages 643–662. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Kawamura:2010:CTO

[KC10] Akitoshi Kawamura and Stephen Cook. Complexity theory for operators in analysis. In ACM [ACM10], pages 495–502. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Keevash:2013:PTP

[KKM13] Peter Keevash, Fiachra Knox, and Richard Mycroft. Polynomialtime perfect matchings in dense hypergraphs. In ACM [ACM13], pages 311–320. ISBN 1-4503-2029-5.

Kaufman:2012:ETR

[KL12] Tali Kaufman and Alexander Lubotzky. Edge transitive Ramanujan graphs and symmetric LDPC good codes. In ACM [ACM12], pages 359–366. ISBN 1-4503-1245-4. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Klauck:2010:SDP

[Kla10] Hartmut Klauck. A strong direct product theorem for disjointness. In ACM [ACM10], pages 77–86. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Kwok:2013:ICI

[KLL^{+}13] Tsz Chiu Kwok, Lap Chi Lau, Yin Tat Lee, Shayan Oveis Gharan, and Luca Trevisan. Improved Cheeger’s inequality: analysis of spectral partitioning algorithms through higher order spectral gap. In ACM [ACM13], pages 11–20. ISBN 1-4503-2029-5.

Kuhn:2010:DCD

[KLO10] Fabian Kuhn, Nancy Lynch, and Rotem Oshman. Distributed computation in dynamic networks. In ACM [ACM10], pages 513– 522. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Kuperberg:2012:PER

[KLP12] Greg Kuperberg, Shachar Lovett, and Ron Peled. Probabilistic existence of rigid combinatorial structures. In ACM [ACM12], pages 1091–1106. ISBN 1-4503-1245-4. LCCN ???? URL http:/

/www.gbv.de/dms/tib-ub-hannover/63314455x..

Kantor:2011:TWC

[KLPP11] Erez Kantor, Zvi Lotker, Merav Parter, and David Peleg. The topology of wireless communication. In ACM [ACM11], pages 383–392. ISBN 1-4503-0691-8. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Khot:2011:NHA

[KM11] Subhash Khot and Dana Moshkovitz. NP-hardness of approximately solving linear equations over reals. In ACM [ACM11], pages 413–420. ISBN 1-4503-0691-8. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Kane:2013:PLF

[KM13] Daniel M. Kane and Raghu Meka. A PRG for Lipschitz functions of polynomials with applications to sparsest cut. In ACM [ACM13], pages 1–10. ISBN 1-4503-2029-5.

Kelner:2012:FAM

[KMP12] Jonathan A. Kelner, Gary L. Miller, and Richard Peng. Faster approximate multicommodity flow using quadratically coupled flows. In ACM [ACM12], pages 1–18. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Klein:2013:SRS

[KMS13] Philip N. Klein, Shay Mozes, and Christian Sommer. Structured recursive separator decompositions for planar graphs in linear time. In ACM [ACM13], pages 505–514. ISBN 1-4503-2029-5.

Kolla:2010:SSN

[KMST10] Alexandra Kolla, Yury Makarychev, Amin Saberi, and Shang-Hua Teng. Subgraph sparsification and nearly optimal ultrasparsifiers. In ACM [ACM10], pages 57–66. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Karnin:2010:DIT

[KMSV10] Zohar S. Karnin, Partha Mukhopadhyay, Amir Shpilka, and Ilya Volkovich. Deterministic identity testing of depth-4 multilinear circuits with bounded top fan-in. In ACM [ACM10], pages 649– 658. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Karande:2011:OBM

[KMT11] Chinmay Karande, Aranyak Mehta, and Pushkar Tripathi. Online bipartite matching with unknown distributions. In ACM [ACM11], pages 587–596. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Kalai:2010:ELM

[KMV10] Adam Tauman Kalai, Ankur Moitra, and Gregory Valiant. Efficiently learning mixtures of two Gaussians. In ACM [ACM10], pages 553–562. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010.

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Koucky:2011:PGG

[KNP11] Michal Koucký, Prajakta Nimbhorkar, and Pavel Pudlák. Pseudorandom generators for group products: extended abstract. In ACM [ACM11], pages 263–272. ISBN 1-4503-0691-8. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Kane:2011:FME

[KNPW11] Daniel M. Kane, Jelani Nelson, Ely Porat, and David P. Woodruff. Fast moment estimation in data streams in optimal space. In ACM [ACM11], pages 745–754. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Kleinberg:2011:MMA

[KO11] Jon Kleinberg and Sigal Oren. Mechanisms for (mis)allocating scientific credit. In ACM [ACM11], pages 529–538. ISBN 1-

4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Kopparty:2011:CPF

[Kop11] Swastik Kopparty. On the complexity of powering in finite fields. In ACM [ACM11], pages 489–498. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Kelner:2013:SCA

[KOSZ13] Jonathan A. Kelner, Lorenzo Orecchia, Aaron Sidford, and Zeyuan Allen Zhu. A simple, combinatorial algorithm for solving SDD systems in nearly-linear time. In ACM [ACM13], pages 911–920. ISBN 1-4503-2029-5.

Khot:2012:LHC

[KPV12] Subhash A. Khot, Preyas Popat, and Nisheeth K. Vishnoi. 2^{log^{1−ϵ} n} hardness for the closest vector problem with preprocessing. In ACM [ACM12], pages 277–288. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Kol:2013:ICC

[KR13a] Gillat Kol and Ran Raz. Interactive channel capacity. In ACM [ACM13], pages 715–724. ISBN 1-4503-2029-5.

Komargodski:2013:ACL

[KR13b] Ilan Komargodski and Ran Raz. Average-case lower bounds for formula size. In ACM [ACM13], pages 171–180. ISBN 1-4503-2029-5.

Kalai:2013:DBS

[KRR13] Yael Tauman Kalai, Ran Raz, and Ron D. Rothblum. Delegation for bounded space. In ACM [ACM13], pages 565–574. ISBN 1-4503-2029-5.

Kasiviswanathan:2010:PPR

[KRSU10] Shiva Prasad Kasiviswanathan, Mark Rudelson, Adam Smith, and Jonathan Ullman. The price of privately releasing contingency tables and the spectra of random matrices with correlated rows. In ACM [ACM10], pages 775–784. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Kopparty:2010:LLD

[KS10] Swastik Kopparty and Shubhangi Saraf. Local list-decoding and testing of random linear codes from high error. In ACM [ACM10], pages 417–426. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010.

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Kolipaka:2011:MTM

[KS11] Kashyap Babu Rao Kolipaka and Mario Szegedy. Moser and Tardos meet Lovász. In ACM [ACM11], pages 235–244. ISBN 1-

4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

King:2013:BAP

[KS13] Valerie King and Jared Saia. Byzantine agreement in polynomial expected time: [extended abstract]. In ACM [ACM13], pages 401– 410. ISBN 1-4503-2029-5.

Kopparty:2011:HRC

[KSY11] Swastik Kopparty, Shubhangi Saraf, and Sergey Yekhanin. Highrate codes with sublinear-time decoding. In ACM [ACM11], pages 167–176. ISBN 1-4503-0691-8. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Kempe:2011:PRE

[KV11] Julia Kempe and Thomas Vidick. Parallel repetition of entangled games. In ACM [ACM11], pages 353–362. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Kleinberg:2012:MPI

[KW12] Robert Kleinberg and Seth Matthew Weinberg. Matroid prophet inequalities. In ACM [ACM12], pages 123–136. ISBN 1-4503-

1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Larsen:2012:CPC

[Lar12] Kasper Green Larsen. The cell probe complexity of dynamic range counting. In ACM [ACM12], pages 85–94. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Lopez-Alt:2012:FMC

[LATV12] Adriana López-Alt, Eran Tromer, and Vinod Vaikuntanathan. On-the-fly multiparty computation on the cloud via multikey fully homomorphic encryption. In ACM [ACM12], pages 1219–1234. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Lee:2012:MWS

[LGT12] James R. Lee, Shayan Oveis Gharan, and Luca Trevisan. Multiway spectral partitioning and higher-order Cheeger inequalities. In ACM [ACM12], pages 1117–1130. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Li:2012:DEN

[Li12] Xin Li. Design extractors, non-malleable condensers and privacy amplification. In ACM [ACM12], pages 837–854. ISBN 1-4503-

1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Li:2013:NIS

[Li13] Xin Li. New independent source extractors with exponential improvement. In ACM [ACM13], pages 783–792. ISBN 1-4503-2029-5.

Lewko:2011:HLK

[LLW11] Allison Lewko, Mark Lewko, and Brent Waters. How to leak on key updates. In ACM [ACM11], pages 725–734. ISBN 1-4503-

0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Lee:2010:BSM

[LM10a] James R. Lee and Mohammad Moharrami. Bilipschitz snowflakes and metrics of negative type. In ACM [ACM10], pages 621–630. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Leighton:2010:ELV

[LM10b] F. Thomson Leighton and Ankur Moitra. Extensions and limits to vertex sparsification. In ACM [ACM10], pages 47–56. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Lee:2013:NCO

[LMM13] James R. Lee, Manor Mendel, and Mohammad Moharrami. A node-capacitated Okamura–Seymour theorem. In ACM [ACM13], pages 495–504. ISBN 1-4503-2029-5.

Lokshtanov:2010:SSA

[LN10] Daniel Lokshtanov and Jesper Nederlof. Saving space by algebraization. In ACM [ACM10], pages 321–330. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Lin:2011:CRN

[LP11] Huijia Lin and Rafael Pass. Constant-round non-malleable commitments from any one-way function. In ACM [ACM11], pages 705–714. ISBN 1-4503-0691-8. LCCN ???? URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Lenzen:2013:FRT

[LPS13] Christoph Lenzen and Boaz Patt-Shamir. Fast routing table construction using small messages: extended abstract. In ACM [ACM13], pages 381–390. ISBN 1-4503-2029-5.

Leonardi:2012:PFA

[LR12] Stefano Leonardi and Tim Roughgarden. Prior-free auctions with ordered bidders. In ACM [ACM12], pages 427–434. ISBN 1-

4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Lee:2013:NAC

[LRS13] Yin Tat Lee, Satish Rao, and Nikhil Srivastava. A new approach to computing maximum flows using electrical flows. In ACM [ACM13], pages 755–764. ISBN 1-4503-2029-5.

Louis:2012:MSC

[LRTV12] Anand Louis, Prasad Raghavendra, Prasad Tetali, and Santosh Vempala. Many sparse cuts via higher eigenvalues. In ACM [ACM12], pages 1131–1140. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Lee:2011:NOD

[LS11] James R. Lee and Anastasios Sidiropoulos. Near-optimal distortion bounds for embedding doubling spaces into L_{1}. In ACM [ACM11], pages 765–772. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Li:2013:AKM

[LS13] Shi Li and Ola Svensson. Approximating k-median via pseudoapproximation. In ACM [ACM13], pages 901–910. ISBN 1-4503-2029-5.

Lee:2010:MMP

[LSV10] Jon Lee, Maxim Sviridenko, and Jan Vondrak. Matroid matching: the power of local search. In ACM [ACM10], pages 369–378. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Lenzen:2011:TBP

[LW11] Christoph Lenzen and Roger Wattenhofer. Tight bounds for parallel randomized load balancing: extended abstract. In ACM [ACM11], pages 11–20. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Li:2013:SCO

[LY13] Jian Li and Wen Yuan. Stochastic combinatorial optimization via Poisson approximation. In ACM [ACM13], pages 971–980. ISBN 1-4503-2029-5.

Madry:2010:FAS

[Mad10] Aleksander Madry. Faster approximation schemes for fractional multicommodity flow problems via dynamic graph algorithms. In ACM [ACM10], pages 121–130. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Marx:2010:THP

[Mar10] Dániel Marx. Tractable hypergraph properties for constraint satisfaction and conjunctive queries. In ACM [ACM10], pages 735–744. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Mendes:2013:MAA

[MH13] Hammurabi Mendes and Maurice Herlihy. Multidimensional approximate agreement in Byzantine asynchronous systems. In ACM [ACM13], pages 391–400. ISBN 1-4503-2029-5.

Miller:2013:SLO

[Mil13] Gary L. Miller. Solving large optimization problems using spectral graph theory. In ACM [ACM13], page 981. ISBN 1-4503-2029-5.

Meng:2013:LDS

[MM13] Xiangrui Meng and Michael W. Mahoney. Low-distortion subspace embeddings in input-sparsity time and applications to robust linear regression. In ACM [ACM13], pages 91–100. ISBN 1-4503-2029-5.

Magniez:2010:RWP

[MMN10] Frédéric Magniez, Claire Mathieu, and Ashwin Nayak. Recognizing well-parenthesized expressions in the streaming model. In ACM [ACM10], pages 261–270. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Makarychev:2012:AAS

[MMV12] Konstantin Makarychev, Yury Makarychev, and Aravindan Vijayaraghavan. Approximation algorithms for semi-random partitioning problems. In ACM [ACM12], pages 367–384. ISBN 1-

4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Muthukrishnan:2012:OPH

[MN12] S. Muthukrishnan and Aleksandar Nikolov. Optimal private halfspace counting via discrepancy. In ACM [ACM12], pages 1285– 1292. ISBN 1-4503-1245-4. LCCN ???? URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Moitra:2011:POS

[MO11] Ankur Moitra and Ryan O’Donnell. Pareto optimal solutions for smoothed analysts. In ACM [ACM11], pages 225–234. ISBN 1-

4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Molloy:2012:FTK

[Mol12] Michael Molloy. The freezing threshold for k-colourings of a random graph. In ACM [ACM12], pages 921–930. ISBN 1-4503-

1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Montanari:2010:MPA

[Mon10] Andrea Montanari. Message passing algorithms: a success looking for theoreticians. In ACM [ACM10], pages 37–38. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

Marx:2011:FPT

[MR11] Dániel Marx and Igor Razgon. Fixed-parameter tractability of multicut parameterized by the size of the cutset. In ACM [ACM11], pages 469–478. ISBN 1-4503-0691-8. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Mossel:2012:QGS

[MR12] Elchanan Mossel and Miklós Z. Rácz. A quantitative Gibbard– Satterthwaite theorem without neutrality. In ACM [ACM12], pages 1041–1060. ISBN 1-4503-1245-4. LCCN ???? URL http:/

/www.gbv.de/dms/tib-ub-hannover/63314455x..

Moser:2011:FDS

[MS11] Robin A. Moser and Dominik Scheder. A full derandomization of Schöning’s k-SAT algorithm. In ACM [ACM11], pages 245–252. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Micciancio:2010:DSE

[MV10] Daniele Micciancio and Panagiotis Voulgaris. A deterministic single exponential time algorithm for most lattice problems based on Voronoi cell computations. In ACM [ACM10], pages 351–358. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Miles:2013:SCG

[MV13] Eric Miles and Emanuele Viola. Shielding circuits with groups. In ACM [ACM13], pages 251–260. ISBN 1-4503-2029-5.

Mahdian:2011:OBM

[MY11] Mohammad Mahdian and Qiqi Yan. Online bipartite matching with random arrivals: an approach based on strongly factorrevealing LPs. In ACM [ACM11], pages 597–606. ISBN 1-4503-

0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Meka:2010:PGP

[MZ10] Raghu Meka and David Zuckerman. Pseudorandom generators for polynomial threshold functions. In ACM [ACM10], pages 427–436. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Nelson:2013:SLB

[NN13] Jelani Nelson and Huy L. Nguyen. Sparsity lower bounds for dimensionality reducing maps. In ACM [ACM13], pages 101–110. ISBN 1-4503-2029-5.

Naor:2013:ERN

[NRV13] Assaf Naor, Oded Regev, and Thomas Vidick. Efficient rounding for the noncommutative Grothendieck inequality. In ACM [ACM13], pages 71–80. ISBN 1-4503-2029-5.

Newman:2011:EPH

[NS11] Ilan Newman and Christian Sohler. Every property of hyperfinite graphs is testable. In ACM [ACM11], pages 675–684. ISBN 1-

4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Neiman:2013:SDA

[NS13] Ofer Neiman and Shay Solomon. Simple deterministic algorithms for fully dynamic maximal matching. In ACM [ACM13], pages 745–754. ISBN 1-4503-2029-5.

Novocin:2011:LRA

[NSV11] Andrew Novocin, Damien Stehlé, and Gilles Villard. An LLLreduction algorithm with quasi-linear time complexity: extended abstract. In ACM [ACM11], pages 403–412. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Nikolov:2013:GDP

[NTZ13] Aleksandar Nikolov, Kunal Talwar, and Li Zhang. The geometry of differential privacy: the sparse and approximate cases. In ACM [ACM13], pages 351–360. ISBN 1-4503-2029-5.

Onak:2010:MLM

[OR10] Krzysztof Onak and Ronitt Rubinfeld. Maintaining a large matching and a small vertex cover. In ACM [ACM10], pages 457–464.

ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Orlin:2010:IAC

[Orl10] James B. Orlin. Improved algorithms for computing Fisher’s market clearing prices. In ACM [ACM10], pages 291–300. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Orlin:2013:MFN

[Orl13] James B. Orlin. Max flows in O(nm) time, or better. In ACM [ACM13], pages 765–774. ISBN 1-4503-2029-5.

Orecchia:2012:AEL

[OSV12] Lorenzo Orecchia, Sushant Sachdeva, and Nisheeth K. Vishnoi. Approximating the exponential, the Lanczos method and an Õ(m)-time spectral algorithm for balanced separator. In ACM [ACM12], pages 1141–1160. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

ODonnell:2012:NPN

[OW12] Ryan O’Donnell and John Wright. A new point of NP-hardness for unique games. In ACM [ACM12], pages 289–306. ISBN 1-

4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Pass:2011:LPS

[Pas11] Rafael Pass. Limits of provable security from standard assumptions. In ACM [ACM11], pages 109–118. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Patrascu:2010:TPL

[Pat10] Mihai Patrascu. Towards polynomial lower bounds for dynamic problems. In ACM [ACM10], pages 603–610. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Paturi:2010:CCS

[PP10] Ramamohan Paturi and Pavel Pudlak. On the complexity of circuit satisfiability. In ACM [ACM10], pages 241–250. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Papadimitriou:2011:OSI

[PP11] Christos H. Papadimitriou and George Pierrakos. On optimal single-item auctions. In ACM [ACM11], pages 119–128. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Patrascu:2011:DRU

[PT11a] Mihai Pătraşcu and Mikkel Thorup. Don’t rush into a union: take time to find your roots. In ACM [ACM11], pages 559–568. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Patrascu:2011:PST

[PT11b] Mihai Patrascu and Mikkel Thorup. The power of simple tabulation hashing. In ACM [ACM11], pages 1–10. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Raz:2010:TRL

[Raz10] Ran Raz. Tensor-rank and lower bounds for arithmetic formulas. In ACM [ACM10], pages 659–666. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Regev:2011:QOW

[RK11] Oded Regev and Bo’az Klartag. Quantum one-way communication can be exponentially stronger than classical communication. In ACM [ACM11], pages 31–40. ISBN 1-4503-0691-8. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Roth:2010:IPM

[RR10] Aaron Roth and Tim Roughgarden. Interactive privacy via the median mechanism. In ACM [ACM10], pages 765–774. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Raghavendra:2010:GEU

[RS10] Prasad Raghavendra and David Steurer. Graph expansion and the unique games conjecture. In ACM [ACM10], pages 755–764. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Raghavendra:2010:AIS

[RST10] Prasad Raghavendra, David Steurer, and Prasad Tetali. Approximations for the isoperimetric and spectral profile of graphs and related parameters. In ACM [ACM10], pages 631–640. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Rothblum:2013:IPP

[RVW13] Guy N. Rothblum, Salil Vadhan, and Avi Wigderson. Interactive proofs of proximity: delegating computation in sublinear time. In ACM [ACM13], pages 793–802. ISBN 1-4503-2029-5.

Roditty:2013:FAA

[RW13] Liam Roditty and Virginia Vassilevska Williams. Fast approximation algorithms for the diameter and radius of sparse graphs. In ACM [ACM13], pages 515–524. ISBN 1-4503-2029-5.

Sharathkumar:2012:NLT

[SA12] R. Sharathkumar and Pankaj K. Agarwal. A near-linear time ϵ-approximation algorithm for geometric bipartite matching. In ACM [ACM12], pages 385–394. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Sherstov:2010:OBS

[She10] Alexander A. Sherstov. Optimal bounds for sign-representing the intersection of two halfspaces by polynomials. In ACM [ACM10], pages 523–532. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010.

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Sherstov:2011:SDP

[She11] Alexander A. Sherstov. Strong direct product theorems for quantum communication and query complexity. In ACM [ACM11], pages 41–50. ISBN 1-4503-0691-8. LCCN ???? URL http://

www.gbv.de/dms/tib-ub-hannover/63314455x..

Sherstov:2012:MPR

[She12a] Alexander A. Sherstov. Making polynomials robust to noise. In ACM [ACM12], pages 747–758. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Sherstov:2012:MCC

[She12b] Alexander A. Sherstov. The multiparty communication complexity of set disjointness. In ACM [ACM12], pages 525–548. ISBN

1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Sherstov:2013:CLB

[She13] Alexander A. Sherstov. Communication lower bounds using directional derivatives. In ACM [ACM13], pages 921–930. ISBN 1-4503-2029-5.

Sarma:2011:DVH

[SHK^{+}11] Atish Das Sarma, Stephan Holzer, Liah Kor, Amos Korman, Danupon Nanongkai, Gopal Pandurangan, David Peleg, and Roger Wattenhofer. Distributed verification and hardness of distributed approximation. In ACM [ACM11], pages 363–372. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Smith:2011:PPS

[Smi11] Adam Smith. Privacy-preserving statistical estimation with optimal convergence rates. In ACM [ACM11], pages 813–822. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Saxena:2011:BIT

[SS11] Nitin Saxena and C. Seshadhri. Blackbox identity testing for bounded top fanin depth-3 circuits: the field doesn’t matter. In ACM [ACM11], pages 431–440. ISBN 1-4503-0691-8. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Sinclair:2013:LYT

[SS13] Alistair Sinclair and Piyush Srivastava. Lee–Yang theorems and the complexity of computing averages. In ACM [ACM13], pages 625–634. ISBN 1-4503-2029-5.

Syrgkanis:2013:CEM

[ST13] Vasilis Syrgkanis and Eva Tardos. Composable and efficient mechanisms. In ACM [ACM13], pages 211–220. ISBN 1-4503-2029-5.

Saraf:2011:BBI

[SV11] Shubhangi Saraf and Ilya Volkovich. Black-box identity testing of depth-4 multilinear circuits. In ACM [ACM11], pages 421–430.

ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Svensson:2010:CHP

[Sve10] Ola Svensson. Conditional hardness of precedence constrained scheduling on identical machines. In ACM [ACM10], pages 745– 754. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Svensson:2011:SCS

[Sve11] Ola Svensson. Santa Claus schedules jobs on unrelated machines. In ACM [ACM11], pages 617–626. ISBN 1-4503-0691-8.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Sohler:2011:SEN

[SW11] Christian Sohler and David P. Woodruff. Subspace embeddings for the L_{1}-norm with applications. In ACM [ACM11], pages 755–764. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Talagrand:2010:MSS

[Tal10] Michel Talagrand. Are many small sets explicitly small? In ACM [ACM10], pages 13–36. ISBN 1-60558-817-2. LCCN QA 76.6

.A152 2010. URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Thorup:2013:BPS

[Tho13] Mikkel Thorup. Bottom-k and priority sampling, set similarity and subset sums with minimal independence. In ACM [ACM13], pages 371–380. ISBN 1-4503-2029-5.

Ta-Shma:2013:IWC

[TS13] Amnon Ta-Shma. Inverting well conditioned matrices in quantum logspace. In ACM [ACM13], pages 881–890. ISBN 1-4503-2029-5.

Thapper:2013:CFV

[TZ13] Johan Thapper and Stanislav Zivny. The complexity of finitevalued CSPs. In ACM [ACM13], pages 695–704. ISBN 1-4503-2029-5.

Ullman:2013:ACQ

[Ull13] Jonathan Ullman. Answering n_{2+o(1)} counting queries with differential privacy is hard. In ACM [ACM13], pages 361–370. ISBN 1-4503-2029-5.

Varadarajan:2010:WGS

[Var10] Kasturi Varadarajan. Weighted geometric set cover via quasiuniform sampling. In ACM [ACM10], pages 641–648. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.gbv.

de/dms/tib-ub-hannover/63314455x..

Vondrak:2011:SFM

[VCZ11] Jan Vondrák, Chandra Chekuri, and Rico Zenklusen. Submodular function maximization via the multilinear relaxation and contention resolution schemes. In ACM [ACM11], pages 783–792. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/

dms/tib-ub-hannover/63314455x..

VAgh:2010:AUN

[Vég10] László A. Végh. Augmenting undirected node-connectivity by one. In ACM [ACM10], pages 563–572. ISBN 1-60558-817-2. LCCN

QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Vegh:2012:SPA

[Vég12] László A. Végh. Strongly polynomial algorithm for a class of minimum-cost flow problems with separable convex objectives. In ACM [ACM12], pages 27–40. ISBN 1-4503-1245-4. LCCN ????

URL http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Valiant:2011:EUS

[VV11] Gregory Valiant and Paul Valiant. Estimating the unseen: an n/ log(n)-sample estimator for entropy and support size, shown optimal via new CLTs. In ACM [ACM11], pages 685–694. ISBN

1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Vazirani:2012:CQD

[VV12] Umesh Vazirani and Thomas Vidick. Certifiable quantum dice: or, true random number generation secure against quantum adversaries. In ACM [ACM12], pages 61–76. ISBN 1-4503-1245-4.

LCCN ???? URL http://www.gbv.de/dms/tib-ub-hannover/
63314455x..

Verbin:2010:LBT

[VZ10] Elad Verbin and Qin Zhang. The limits of buffering: a tight lower bound for dynamic membership in the external memory model. In ACM [ACM10], pages 447–456. ISBN 1-60558-817-2.

LCCN QA 76.6 .A152 2010. URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Vadhan:2012:CPS

[VZ12] Salil Vadhan and Colin Jia Zheng. Characterizing pseudoentropy and simplifying pseudorandom generator constructions. In ACM [ACM12], pages 817–836. ISBN 1-4503-1245-4. LCCN ???? URL

http://www.gbv.de/dms/tib-ub-hannover/63314455x..

Williams:2010:IES

[Wil10] Ryan Williams. Improving exhaustive search implies superpolynomial lower bounds. In ACM [ACM10], pages 231–240. ISBN 1-60558-817-2. LCCN QA 76.6 .A152 2010. URL http://www.

gbv.de/dms/tib-ub-hannover/63314455x..

Williams:2012:MMF

[Wil12] Virginia Vassilevska Williams. Multiplying matrices faster than Coppersmith–Winograd. In ACM [ACM12], pages 887–898. ISBN

1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Williams:2013:NPV

[Wil13] Ryan Williams. Natural proofs versus derandomization. In ACM [ACM13], pages 21–30. ISBN 1-4503-2029-5.

Woodruff:2011:NOP

[Woo11] David P. Woodruff. Near-optimal private approximation protocols via a black box transformation. In ACM [ACM11], pages 735–744. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..

Wootters:2013:LDR

[Woo13] Mary Wootters. On the list decodability of random linear codes with large error rates. In ACM [ACM13], pages 853–860. ISBN 1-4503-2029-5.

Woodruff:2012:TBD

[WZ12] David P. Woodruff and Qin Zhang. Tight bounds for distributed functional monitoring. In ACM [ACM12], pages 941–960. ISBN

1-4503-1245-4. LCCN ???? URL http://www.gbv.de/dms/tib-
ub-hannover/63314455x..

Yoshida:2011:OCT

[Yos11] Yuichi Yoshida. Optimal constant-time approximation algorithms and (unconditional) inapproximability results for every boundeddegree CSP. In ACM [ACM11], pages 665–674. ISBN 1-4503-

0691-8. LCCN ???? URL http://www.gbv.de/dms/tib-ub-
hannover/63314455x..

Zewi:2011:ATS

[ZBS11] Noga Zewi and Eli Ben-Sasson. From affine to two-source extractors via approximate duality. In ACM [ACM11], pages 177–186. ISBN 1-4503-0691-8. LCCN ???? URL http://www.gbv.de/dms/

tib-ub-hannover/63314455x..